### v1

# Geometric Hardy inequalities on the Heisenberg groups via convexity

Abstract.
We prove $L^{p}$-Hardy inequalities with distance to the boundary for domains in the Heisenberg group ${\mathbb{H}}^{n}$, $n\geq 1$. Our results are based on a geometric condition. This is first implemented for the Euclidean distance in certain non-convex domains. It is also implemented on half-spaces and convex polytopes for the distance defined by the gauge quasi-norm on ${\mathbb{H}}^{n}$ related to the fundamental solution of the horizontal Laplacian. In the more general context of a stratified Lie group of step two we study the superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an alternative proof for the $L^{2}$-Hardy inequality on convex domains. In all cases the constants are shown to be sharp.
Key words and phrases:
Geometric $L^{p}$-Hardy inequalities; gauge pseudodistance; Heisenberg group; stratified groups of step two; horizontal convexity
2020 Mathematics Subject Classification:
Primary 26D10, 35R03; Secondary 35A23, 35H20, 35J75
M. Chatzakou is supported by the FWO Odysseus 1 grant G.0H94.18N: Analysis and Partial Differential Equations and by the Methusalem programme of the Ghent University Special Research Fund (BOF) (Grant number 01M01021), and is a postdoctoral fellow of the Research Foundation-Flanders (FWO) under the postdoctoral grant No 12B1223N.
1. Introduction
The classical $L^{p}$-Hardy inequality, $p>1$, affirms that
$\int_{\mathbb{R}^{n}}|\nabla u|^{p}\,{\rm d}x\geq\left|\frac{n-p}{p}\right|^{p}\int_{\mathbb{R}^{n}}\frac{|u|^{p}}{|x|^{p}}\,{\rm d}x\,,$
for $u\in C_{c}^{\infty}(\mathbb{R}^{n}\setminus\{0\})$, where the constant is sharp.
Another much studied type of Hardy inequality is where the  Hardy potential is the distance to the boundary of a reference domain. A well known such result states that if $\Omega\subset\mathbb{R}^{n}$ is a convex domain and $d(x)={\rm dist}(x,\partial\Omega)$, then for any $u\in{C^{\infty}_{c}}(\Omega)$ there holds
$\int_{\Omega}|\nabla u|^{p}\,{{\rm d}}x\geq\left(\frac{p-1}{p}\right)^{p}\int_{\Omega}\frac{|u|^{p}}{d^{p}}\,{{\rm d}}x\,,$
and the constant is sharp, cf. [MS97]. In [BFT04] the convexity condition was replaced by the more general notion of weak mean convexity, namely the requirement that $\Delta d\leq 0$ in the distributional sense in $\Omega$. The above inequality is not valid without some geometric assumptions on $\Omega$ and for this reason inequalities of this type are often called geometric Hardy inequalities. The literature on geometric Hardy inequalities in Euclidean space is large and we refer the interested reader to the works [Ba24, BEL15, Dav98, RS19] which provide an overview of the topic.
On the other hand, subelliptic Hardy inequalities have been studied for quite a long time and the work [GL90] of Garofallo and Lanconelli in the 90’s opened up the research in this direction. By subelliptic Hardy inequalities, we mean Hardy-type inequalities considered in the setting of homogeneous Lie groups, and in particular stratified groups. The systematic analysis of homogeneous Lie groups goes back to the seminal work [FS82] by Folland and Stein where the authors establish the corresponding “anisotropic” non-commutative harmonic analysis, see also [S93]. In view of their importance in the area of partial differential equations, stratified Lie groups, have been widely recognised as they play a key role in establishing subelliptic estimates for differential operators on general manifolds.
As in the Euclidean case, Hardy inequalities on stratified groups may involve either the distance to a point or the distance to the boundary. Moreover, one may use the Euclidean distance, the Carnot-Carathéodory distance or the distance related to the fundamental solution of the sub-Laplacian $\Delta_{H}$, often called the gauge pseudodistance.
Concerning the distance to a point case we refer to [CCR15, D’A04, GL90, GKY17, FP21, RS17, Y13]. For an overview of the works in Hardy inequalities of all the above types we refer to the monograph [RS19]; see also the survey article [Su22].
For Hardy inequalities involving the distance to the boundary the literature in the stratified setting is limited and in most cases it involves the Euclidean distance. In [Lar16] ${\mathbb{H}}^{n}$, S. Larson shows that if $\Omega\subset{\mathbb{H}}^{n}$ is either a half space or a convex (in the Euclidean sense) domain, then for any $u\in C_{c}^{\infty}(\Omega)$ we have
$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{2}\,{\rm d}x\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{2}}{d^{2}}u^{2}\,{\rm d}x\,,$
where $d(x)$, $x\in\Omega$, stands for the Euclidean distance to $\partial\Omega$ and the constant $1/4$ is the best possible. Later on in [RSS20] Ruzhansky et al. proved that if $\Omega$ is a half-space in any stratified group then for any $p>1$ there holds
$\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}x\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{T}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
where $\nabla_{H}$ denotes the horizontal gradient on $H$. For other results in this direction see also [Rus18].
Our main interest in this work is to prove subelliptic geometric Hardy inequalities with best constant in the Heisenberg group ${\mathbb{H}}^{n}$ but also on any stratified group of step two. Our approach is based on the general method of [BFT04] and in particular in the $L^{p}$-superharmonicity of the distance function. The general result is then implemented in different contexts.
In the case of the Euclidean distance on the Heisenberg group ${\mathbb{H}}^{n}$ we prove the following result which goes beyond the convexity condition. The precise value of the constant $\beta(p,n)$ is given in Proposition 3.
Theorem A. Let $R>\rho>0$ and let $T$ denote the torus
$T=\{\xi=(x,y,t)\in{\mathbb{H}}^{n}:\;(r-R)^{2}+t^{2}<\rho^{2}\}$
where $r=\sqrt{|x|^{2}+|y|^{2}}$. For any $p>1$ there exists a positive constant $\beta(p,n)$ such that if
$({\rm i})\qquad$ $R\geq\rho+\Big{(}\frac{(2n-1)\rho}{4}\Big{)}^{\frac{1}{3}}$
$({\rm ii})\qquad$ $R\geq\beta(p,n)\rho,$
then
$\int_{T}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{T}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{p}}{d^{p}}|u|^{p}{\rm d}\xi\;,\qquad u\in{C^{\infty}_{c}}(T).$
Another class of distances in a stratified group consists of those induced by any homogenous quasi-norm. In the case of the Heisenberg group $\mathbb{H}^{n}$ one important such quasi-norm is
(1) $N(\xi)=\Big{(}(|x|^{2}+|y|^{2})^{2}+t^{2}\Big{)}^{\frac{1}{4}},\qquad\xi=(x,y,t)\in{\mathbb{H}}^{n}.$
We recall if we set $Q=2n+2$,the homogeneous dimension of $\mathbb{H}^{n}$, then up to a multiplicative constant, $N^{2-Q}$ is the fundamental solution of the sub-Laplacian on the Heisenberg group ${\mathbb{H}}^{n}$ [Fol73]. Denoting by $d_{N}$ the distance to the boundary incuded by $N$, cf. (12) below, we have the following
Theorem B. (i) Let $p>1$ and let $D\subset{\mathbb{H}}^{n}$ be a half-space. There holds
$\int_{D}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\left(\frac{p-1}{p}\right)^{p}\int_{D}\frac{|\nabla_{{\mathbb{H}}^{n}}d_{N}|^{p}}{d_{N}^{p}}|u|^{p}\,{\rm d}\xi\;,\qquad u\in{C^{\infty}_{c}}(D)\,.$
(ii) In case $p=2$ the above inequality is also valid for any bounded convex polytope.
The constant is the best possible in both cases.
In Section 4 we extend our setting to that of an arbitrary stratified group of step two. In this setting a certain notion of convexity plays a central role. There are various notions of convexity in the sub-Riemannian setting and their properties can vary significantly [DGN03, DLZ24, LMS03]. Notably, in [MR03] R. Monty and M. Rickly prove that in the case of the Heisenberg group $\mathbb{H}$, if a set is geodesically convex and contains at least three points that do not lie on the same geodesic, then it necessarily coincides with $\mathbb{H}$.
In our context the relevant notions of convexity of sets and functions are the ones introduced at the same time by Lu, Manfredi and Stroffolini in [LMS03] on the Heisenberg group and by Danielli, Garofallo and Nhieu in [DGN03] on any stratified group. These notions are the analogues of the corresponding ones in the abelian case $\mathbb{R}^{n}$ but with a twist; the condition refers to a convex combination of two elements $g,g^{\prime}$ for which, additionally, $g^{\prime}\in H_{g}$, i.e. $g^{\prime}$ lies in the horizontal plane passing through $g$; see Section 4.1 for the precise definitions. Exploring properties of the so-called weakly $H$-concave functions (see Definition 12) and using a result from [DGN03] we prove the following theorem, part (iii) of which is contained in [RSS20].
Theorem C. Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
[…]
${\rm(ii)}$ $\quad\Delta_{{\mathbb{H}}^{n}}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
${\rm(iii)}$ The Hardy inequality
$\hskip 42.67912pt\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}g\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{2}}{d^{2}}u^{2}{\rm d}g\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
is valid.
2. two general results on stratified groups
A stratified (or Carnot) group $G\equiv\mathbb{R}^{n}$ is naturally a homogeneous Lie group. Denoting by $\mathfrak{g}$ the corresponding Lie algebra we have $\dim(\mathfrak{g})=n$ and $\mathfrak{g}$ admits a vector space decomposition of the form
(2) $\mathfrak{g}=\bigoplus_{j=1}^{r}V_{j}\,,\quad\text{such that}\quad\left\{\begin{array}[]{l}[V_{1},V_{i-1}]=V_{i}\,,\quad 2\leq i\leq r,\\ [V_{1},V_{r}]=\{0\},\end{array}\right.$
where
$[V_{i},V_{j}]={\rm span}\{[X,Y]:\;X\in V_{i},\;Y\in V_{j}\}.$
Such a stratification naturally equips $G$ with a non-anisotropic dilation structure $\delta_{\lambda}:G\rightarrow G$, $\lambda>0$, and makes $G$ a homogeneous Lie group. The vector spaces $V_{i}$ are called the strata of the Lie algebra $\mathfrak{g}$. A symmetric homogeneous (quasi-)norm on $G$ is a function $N:G\rightarrow[0,\infty)$ such that (i) $N(g)=0$ if and only if $g=e$, where $e$ is the identity element of $G$; (ii) $N(g)=N(g^{-1})$; and (iii) $N(\delta_{\lambda}(g))=\lambda N(g)$. In this article we shall use the term quasi-norm to indicate a symmetric homogeneous quasi-norm.
If $G$ is a stratified group, the system $\{X_{1},\ldots,X_{m}\}$, $m\leq n$, of vector fields in the first stratum $V_{1}$ of $\mathfrak{g}$ generates, after iterated commutators, the whole of $\mathfrak{g}$, and so it is a system of Hörmander vector fields on $\mathbb{R}^{n}$. The vector space spanned by $\{X_{1},\ldots,X_{m}\}$ is referred to as the horizontal hyperplane.
The first-order vector-valued differential operator
$\nabla_{H}=(X_{1},\ldots,X_{m})$
is then called the horizontal gradient on $G$ (or the subgradient on $G$). Similarly ${\rm div}_{H}$ will denote the horizontal divergence given by
${\rm div}_{H}(f_{1},\ldots,f_{m})=X_{1}f_{1}+\ldots X_{m}f_{m}.$
The second-order differential operator
$\Delta_{H}=X_{1}^{2}+\ldots+X_{m}^{2}$
is called the horizontal Laplacian (or sublaplacian) on $G$ and is the sub- Riemannian analogue of the Laplacian on $\mathbb{R}^{n}$. By Hörmander’s Theorem, see [Hör67], the operator $\Delta_{H}$ is hypoelliptic. For $p>1$ we also have the associated horizontal $p$-Laplacian given by
$\Delta_{p,H}\,u={\rm div}_{H}\big{(}|\nabla_{H}u|^{p-2}\nabla_{H}u\big{)}.$
Finally, let us also recall that the (bi-invariant) Haar measure in the case of a stratified group is just, up to multiplication by a constant, the Lebesgue measure on the underlying manifold $\mathbb{R}^{n}$.
We first prove a general theorem which will be later applied in the case of the Euclidean distance and of the pseudodistance induced by the quasi-norm (1).
In what follows, we will say that a function is CC-Lipschitz if it is Lipschitz with respect to the Carnot-Carathéodory distance (equivalently, with respect to the distance induced by any homogeneous quasi-norm).
Part (a) of the next theorem is essentially contained in [RSS20] but we include the short proof of it because of its central role in the present article.
Theorem 1.
Let $G$ be a stratified group and let $\Omega\subset G$ be open and connected. Let $p>1$ and let $d:\Omega\rightarrow(0,\infty)$ be a positive, locally CC-Lipschitz function.
(a) Assume that
$\Delta_{p,H}\,d\leq 0\,\quad\text{in}\,\,\Omega\,,$
where the inequality is understood in the distributional sense. Then
$\int_{\Omega}|\nabla_{H}u|^{p}{{\rm d}x}\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}\,{{\rm d}x}\,,\qquad u\in C_{c}^{\infty}(\Omega).$
(b) Assume that there exist $x_{0}\in\partial\Omega$ and two neighbourhoods $A,A^{\prime}$ of $x_{0}$ in $G$ with $A^{\prime}\subset\subset A$ and such that
${\rm(i)}$ There exists $c>0$ such that $|\nabla_{H}d|\geq c$ in $A\cap\Omega$.
[…]
Then
$\inf_{u\in{C^{\infty}_{c}}(\Omega)}\frac{\int_{\Omega}|\nabla_{H}u|^{p}{{\rm d}x}}{\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{{\rm d}x}}\leq\Big{(}\frac{p-1}{p}\Big{)}^{p}.$
Proof.
First we note that in view of [MSC01, Theorem 2.5] $\nabla_{H}d$ exists a.e. in $\Omega$. Let $T$ be a vector field in $L^{1}_{\rm loc}(\Omega)$ and $u\in C^{\infty}_{c}(\Omega)$. Using an argument from [BFT04], with the only difference that differential operators are replaced by the corresponding horizontal ones, we obtain that
$\int_{\Omega}|\nabla_{H}u|^{p}{{\rm d}x}\geq\int_{\Omega}\Big{(}{\rm div}_{H}T-(p-1)|T|^{\frac{p}{p-1}}\Big{)}|u|^{p}{{\rm d}x},$
where ${\rm div}_{H}T$ is understood in the distributional sense. We now make the particular choice
$T=-\Big{(}\frac{p-1}{p}\Big{)}^{p-1}\frac{1}{d^{p-1}}|\nabla_{H}d|^{p-2}\nabla_{H}d.$
For this choice we have
${\rm div}_{H}T-(p-1)|T|^{\frac{p}{p-1}}=\Big{(}\frac{p-1}{p}\Big{)}^{p}\frac{|\nabla_{H}d|^{p}}{d^{p}}-\Big{(}\frac{p-1}{p}\Big{)}^{p-1}\,\frac{1}{d^{p}}\,\Delta_{p,H}d\,,$
hence (a) follows.
To prove (b), let $\psi$ be a smooth cut-off function supported in $A$ and satisfying $0\leq\psi\leq 1$ and $\psi(x)=1$ in $A^{\prime}$. We fix $\epsilon>0$, which will eventually tend to zero, and we define
$u_{\epsilon}(x)=d(x)^{\frac{p-1}{p}+\epsilon}\psi(x),\qquad x\in\Omega.$
A standard argument shows that $u_{\epsilon}$ can be used as a test function. Applying the elementary inequality $|a+b|^{p}\leq|a|^{p}+c_{p}(|a|^{p-1}|b|+|b|^{p})$, $a,b\in\mathbb{R}^{n}$, we obtain
$|\nabla u_{\epsilon}|^{p}$ $=\Big{|}\big{(}\frac{p-1}{p}+\epsilon\big{)}d^{-\frac{1}{p}+\epsilon}\psi\nabla_{H}d+d^{\frac{p-1}{p}+\epsilon}\nabla_{H}\psi\Big{|}^{p}$
$\leq\!\Big{(}\frac{p-1}{p}+\epsilon\!\Big{)}^{p}\!d^{-1+\epsilon p}\psi^{p}|\nabla_{H}d|^{p}\!+c_{p}^{\prime}d^{\epsilon p}|\nabla_{H}d|^{p-1}|\nabla_{H}\psi|+c_{p}d^{p-1+\epsilon p}|\nabla_{H}\psi|^{p}.$
It follows that
$\int_{\Omega}|\nabla_{H}u_{\epsilon}|^{p}{{\rm d}x}$ $\leq\Big{(}\frac{p-1}{p}+\epsilon\Big{)}^{p}\int_{\Omega}d^{-1+\epsilon p}\psi^{p}|\nabla_{H}d|^{p}{{\rm d}x}$
$\quad+c_{p}^{\prime}\int_{\Omega}d^{\epsilon p}|\nabla_{H}d|^{p-1}|\nabla_{H}\psi|\,{{\rm d}x}+c_{p}\int_{\Omega}d^{p-1+\epsilon p}|\nabla_{H}\psi|^{p}{{\rm d}x}.$
The last two integrals stay bounded as $\epsilon\to 0$, so
$\int_{\Omega}|\nabla_{H}u_{\epsilon}|^{p}{{\rm d}x}\leq\Big{(}\frac{p-1}{p}+\epsilon\Big{)}^{p}\int_{\Omega}d^{-1+\epsilon p}\psi^{p}|\nabla_{H}d|^{p}{{\rm d}x}+O(1).$
We also have
$\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}u_{\epsilon}^{p}{{\rm d}x}=\int_{\Omega}|\nabla_{H}d|^{p}d^{-1+\epsilon p}\psi^{p}{{\rm d}x}\,.$
Hence, since the last integral diverges to infinity as $\epsilon\to 0$, we arrive at
$\frac{\int_{\Omega}|\nabla_{H}u_{\epsilon}|^{p}{{\rm d}x}}{\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}u_{\epsilon}^{p}{{\rm d}x}}=\Big{(}\frac{p-1}{p}+\epsilon\Big{)}^{p}+o(1),\qquad\mbox{ as }\epsilon\to 0.$
Letting $\epsilon\to 0$ concludes the proof. ∎
Given a quasi-norm $N$ we may define a pseudodistance by
$d_{N}(x,y)=N(y^{-1}x)\;,\qquad x,y\in G.$
The following proposition allows us to apply Theorem 1 in the case of the gauge quasi-norm (1).
Proposition 2.
Let $S$ be a closed set in a stratified group $G$ and let $N$ be any quasi-norm on $G$ that is smooth out of the origin. Then the pseudodistance
$d_{N,S}(g)=\inf_{b\in S}d_{N}(b,g)=\inf_{b\in S}N(g^{-1}b)$
is CC-Lipschitz.
Proof.
By the equivalence of all quasi-norms on a stratified group it is enough to show that for $g,g^{\prime}\in S$ we have
(3) $|d(g)-d(g^{\prime})|\leq Kd_{N}(g^{-1}g^{\prime})\,,$
for some $K>0$. By [BLU07, Proposition 5.14.1] there exists $\beta\geq 1$ such that
$N(xy)\leq\beta N(x)+N(y)\,,\quad\mbox{ for all}\,\,x,y\in G\,.$
Now, let $(b_{n}),(b_{n})^{\prime}\subset S$ be such that
$d(g)=\lim N(g^{-1}b_{n})\,,\qquad\quad d(g^{\prime})=\lim N(g^{\prime-1}b_{n}^{\prime})\,,$
We then have
$d(g)-d(g^{\prime})$$\leq$ lim $\liminf\big{[}N(g^{-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
$=$ lim $\liminf\big{[}N(g^{-1}g^{\prime}g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
$\leq$ lim $\liminf\big{[}\beta N(g^{-1}g^{\prime})+N(g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
$=$ $\beta N(g^{-1}g^{\prime})\,.$
Similarly we can show that
$d(g^{\prime})-d(g)\leq C\beta\,d_{N}g^{\prime},g)\,,$
and (3) follows. ∎
3. Geometric Hardy inequalities on the Heisenberg group
In the section we consider geometric Hardy inequalities on the Heisenberg group $\mathbb{H}^{n}$. In the first part we consider the Euclidean distance and prove the validity of the Hardy inequality with best constant on certain torii under suitable assumptions on the radii. In the second part we study the geometric Hardy inequality for the pseudodistance induced by the quasi-norm (1).
We recall that the Heisenberg group $\mathbb{H}^{n}$ is the manifold
$\mathbb{H}^{n}=\{\xi=(x,y,t):x,y\in\mathbb{R}^{n},\;t\in\mathbb{R}\}$
equipped with the group operation
$\xi\xi^{\prime}=\big{(}x+x^{\prime},y+y^{\prime},t+t^{\prime}+2(x\cdot y^{\prime}-y\cdot x^{\prime})\big{)}.$
The left-invariant vector fields
$X_{i}=\partial_{x_{i}}+2y_{i}\partial_{t}\;,\qquad Y_{i}=\partial_{y_{i}}-2x_{i}\partial_{t}\,,\quad i=1,\ldots,n,$
form the canonical basis basis of the first stratum and the associated horizontal gradient and horizontal Laplacian on ${\mathbb{H}}^{n}$ are given respectively by
$\nabla_{{\mathbb{H}}^{n}}=(X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}),$
and
$\Delta_{{\mathbb{H}}^{n}}=\sum_{i=1}^{n}(X_{i}^{2}+Y_{i}^{2})\,.$
So in the current section we denote the horizontal gradient and Laplacian by $\nabla_{{\mathbb{H}}^{n}}$ and $\Delta_{{\mathbb{H}}^{n}}$, respectively, to emphasize that the obtained results refer to the particular case of ${\mathbb{H}}^{n}$.
3.1. Hardy inequalities with respect to the Euclidean distance on a torus
We will see here that Theorem 1 can be applied in the case of the Euclidean distance on a torus and thus goes beyond the convexity assumption of [Lar16].
We first note that for any $u\in C^{2}({\mathbb{H}}^{n})$ we have
(4) $\Delta_{{\mathbb{H}}^{n}}u=\sum_{i=1}^{n}(u_{x_{i}x_{i}}+u_{y_{i}y_{i}})+4(|x|^{2}+|y|^{2})u_{tt}-2\sum_{i=1}^{n}(y_{i}u_{x_{i}t}-x_{i}u_{y_{i}t})$
We now use cylindrical coordinates $(r,\omega,t)$ in ${\mathbb{H}}^{n}$, that is spherical coordinates $(r,\omega)$ in $\mathbb{R}^{2n}$,
$(x,y)=r\omega\;,\quad r>0,\;\omega\in\mathrm{S}^{2n-1},$
where $\mathrm{S}^{2n-1}$ denotes the unit sphere in $\mathbb{R}^{2n}$. The Euclidean gradient in $\mathbb{R}^{2n}$ is then given by
$\nabla u=u_{r}\,\omega+\frac{1}{r}\nabla_{\omega}u.$
Suppose now that a function $u\in C^{2}({\mathbb{H}}^{n})$ is independent of $\omega$, that is $u=u(r,t)$. In this case $\nabla u=u_{r}\,\omega$, so
$\sum_{i=1}^{n}(y_{i}u_{x_{i}}-x_{i}u_{y_{i}})=\sum_{i=1}^{n}(y_{i}\frac{u_{r}}{r}x_{i}-x_{i}\frac{u_{r}}{r}y_{i})=0.$
Hence, for such functions, (4) gives
$\Delta_{{\mathbb{H}}^{n}}u=u_{rr}+\frac{2n-1}{r}u_{r}+4r^{2}u_{tt}.$
Consider now a torus $T\subset\mathbb{H}^{n}$ which is symmetric with respect to the $t$-axis and is centered at the origin. Letting $R,\rho$ ($R>\rho$) denote the two radii, $T$ is described in cylindrical coordinates as
(5) $T=\{\xi=(r,\omega,t):\;(r-R)^{2}+t^{2}<\rho^{2}\}.$
The Euclidean distance to the boundary is given by
$d(\xi)=\rho-\sqrt{(r-R)^{2}+t^{2}},\qquad\xi\in T,$
and is smooth in $T$ except on the $(2n-1)$-dimensional ‘circle’
$S=\{\xi=(r,\omega,t):\;t=0,\;r=R\}.$
The horizontal Laplacian of $d$ is then given by
(6) $\Delta_{\mathbb{H}^{n}}d=d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt},\qquad\mbox{ in }T\setminus S.$
In $T\setminus S$ we have
$d_{r}$ $=-\frac{r-R}{\sqrt{(r-R)^{2}+t^{2}}}$
$d_{rr}$ $=-\big{(}(r-R)^{2}+t^{2}\big{)}^{-\frac{3}{2}}t^{2}$
$d_{tt}$ $=-\big{(}(r-R)^{2}+t^{2}\big{)}^{-\frac{3}{2}}(r-R)^{2}.$
Substituting in (6) we conclude that in $T\setminus S$ there holds
$\Delta_{\mathbb{H}^{n}}d=$ $-\big{(}(r-R)^{2}+t^{2}\big{)}^{-\frac{3}{2}}\frac{1}{r}$
(7) ×$\times\bigg{\{}\Big{[}(2n-1)(r-R)+4r^{3}\Big{]}(r-R)^{2}+\Big{[}r+(2n-1)(r-R)\Big{]}t^{2}\bigg{\}}.$
Proposition 3.
Let $p>1$. Let $R>\rho>0$ and let $T$ be the torus (5). Then there exists a positive constant $\beta(p,n)$ such that if
$({\rm i})\qquad$ $R\geq\rho+\Big{(}\frac{(2n-1)\rho}{4}\Big{)}^{\frac{1}{3}}$
$({\rm ii})\qquad$ $R\geq\beta(p,n)\rho,$
then $\Delta_{p,\mathbb{H}^{n}}d\leq 0$ in $T\setminus S$. Moreover we can take
$\beta(p,n)=\left\{\begin{array}[]{ll}\max\Big{\{}\frac{2n+p-2}{p-1}\,,\;\frac{2n-p+1}{2(2-p)}\Big{\}},&\mbox{ if }1<p<2,\\ 2n&\mbox{ if }p=2,\\ 2n+p-2&\mbox{ if }p\geq\frac{13+\sqrt{32n-7}}{8}.\end{array}\right.$
whereas for $2<p<\frac{13+\sqrt{32n-7}}{8}$ we have $\beta(p,n)=1+\frac{1}{a(p,n)}$, where $a(p,n)$ is the positive solution of
$(2n+p-3)^{2}a^{2}+4\Big{(}(p-2)(2n+p-3)+(p-1)(2n-1)\Big{)}a-4(2p-3)=0$
Proof.
Let $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$. We then have
(8) $\Delta_{p,\mathbb{H}^{n}}d=A^{\frac{p-4}{2}}\Big{(}A\,\Delta_{\mathbb{H}^{n}}d+\frac{p-2}{2}(d_{r}A_{r}+4r^{2}d_{t}A_{t})\Big{)}.$
Now, simple computations give
(9) $A=d_{r}^{2}+4r^{2}d_{t}^{2}=\frac{(r-R)^{2}+4r^{2}t^{2}}{(r-R)^{2}+t^{2}}.$
and
$A_{r}=\frac{2(r-R)(1-4rR)t^{2}+8rt^{4}}{\big{(}(r-R)^{2}+t^{2}\big{)}^{2}},\qquad A_{t}=\frac{2t(r-R)^{2}(4r^{2}-1)}{\big{(}(r-R)^{2}+t^{2}\big{)}^{2}}.$
Hence
$d_{r}A_{r}+4r^{2}d_{t}A_{t}=-\Big{(}(r-R)^{2}+t^{2}\Big{)}^{-\frac{5}{2}}\bigg{\{}(r-R)^{2}t^{2}(2-8r^{2}-8rR+32r^{4})+8r(r-R)t^{4}\bigg{\}}.$
Substituting in (8) and recalling (7) we obtain after some more computations that
$\Delta_{p,\mathbb{H}^{n}}d=-\,A^{\frac{p-4}{2}}\big{(}(r-R)^{2}+t^{2}\big{)}^{-\frac{5}{2}}W\,,$
where
$W$ $=\frac{1}{r}\Big{(}(r-R)^{2}+4r^{2}t^{2}\Big{)}\bigg{(}(r-R)^{2}\Big{[}(2n-1)(r-R)+4r^{3}\Big{]}$
$\quad\hskip 99.58464pt+\Big{[}(2n-1)(r-R)+r\Big{]}t^{2}\bigg{)}$
$\qquad+(p-2)\bigg{(}(r-R)^{2}\Big{[}1-4rR-4r^{2}+16r^{4}\Big{]}t^{2}+4r(r-R)t^{4}\bigg{)}.$
In case $p=2$ we note that our assumption implies that
$(2n-1)(r-R)+4r^{3}\geq 0\;,\quad(2n-1)(r-R)+r\geq 0$
in $T$, hence $W\geq 0$, as required.
For $p\neq 2$ we collect similar powers of $t$ to obtain
$rW$ $=(r-R)^{4}\Big{(}(2n-1)(r-R)+4r^{3}\Big{)}$
$\quad+(r-R)^{2}\bigg{\{}\Big{(}2n-1+4(2n+p-3)r^{2}\Big{)}(r-R)$
$\quad+16(p-1)r^{5}-8(p-2)r^{3}+(p-1)r\bigg{\}}t^{2}$
$\quad+4r^{2}\Big{(}(2n+p-3)(r-R)+r\Big{)}t^{4}$
$=:C_{0}+C_{1}t^{2}+C_{2}t^{4}.$
Assumption (i) implies that $C_{0}\geq 0$. Similarly, assumption (ii) implies that $R\geq(2n+p-2)\rho$ in all cases and therefore $C_{2}\geq 0$.
We shall prove that $C_{1}\geq 0$. Equivalently, that
$r-R+r\,\frac{16(p-1)r^{4}-8(p-2)r^{2}+p-1}{2n-1+4(2n+p-3)r^{2}}\geq 0,\quad\mbox{ for all }R-\rho\leq r\leq R+\rho.$
For this we shall find a positive constant $a=a(p,n)$ such that
$\frac{16(p-1)r^{4}-8(p-2)r^{2}+p-1}{2n-1+4(2n+p-3)r^{2}}\geq a\;,\quad r>0,$
or equivalently
(10) $016(p-1)r^{4}-\Big{(}8(p-2)+4a(2n+p-3)\Big{)}r^{2}+p-1-a(2n-1)\geq 0\,,\;\quad r>$ 0 .
If such an $a$ has been found then we shall have $C_{1}\geq 0$ provided the radii of $T$ satisfy
$r-R+ar\geq 0,\quad\mbox{ for all }r\in[R-\rho,R+\rho],$
which is equivalent to
$R\geq(1+\frac{1}{a})\rho.$
At this point we need to distinguish different cases.
Case $21<p<$ 2. In this case we choose
$a=a_{1}(p,n):=\min\Big{\{}\frac{p-1}{2n-1},\;\frac{2(2-p)}{2n+p-3}\Big{\}}$
which makes all coefficients in (10) non-negative. The requirement on the radii then is
$R\geq\Big{(}1+\frac{1}{a_{1}(p,n)}\Big{)}\rho=\max\Big{\{}\frac{2n+p-2}{p-1}\,,\;\frac{2n-p+1}{2(2-p)}\Big{\}}\,\rho,$
and it is satisfied by our assumptions.
Case $p>2$. In this case the coefficient of $r^{2}$ in (10) is negative, so we consider the discriminant. We have
$\Big{(}8(p-2)+4a(2n+p-3)\Big{)}^{2}-64(p-1)\Big{(}p-1-a(2n-1)\Big{)}$
$=$ $16\bigg{\{}-4(2p-3)+4\Big{(}(p-2)(2n+p-3)+(p-1)(2n-1)\Big{)}a+(2n+p-3)^{2}a^{2}\bigg{\}}.$
We now choose $a=a(p,n)$ to be the positive root of the quadratic polynomial above. So the requirement on the radii for (10) is
$R\geq\Big{(}1+\frac{1}{a(p,n)}\Big{)}\rho$
and $\beta(p,n)$ is given by
[…]
Finally we to note that for $p>2$,
$2n+p-2\geq 1+\frac{1}{a(p,n)}$ $\;\Longleftrightarrow\;a(p,n)\geq\frac{1}{2n+p-3}$
 $\;\Longleftrightarrow\;4p^{2}-13p+11-2n\geq 0$
$\;\Longleftrightarrow\;p\geq\frac{13+\sqrt{32n-7}}{8}.$
This completes the proof. ∎
Remark 1.
For $21<p<$ 2 we have that
$\frac{2n-p+1}{2(2-p)}\geq 2n+p-2\quad\mbox{ iff }\quad\frac{3}{2}\leq p<2.$
Moreover if we define $p_{0}=\sqrt{9n^{2}-8n+2}-3(n-1)$ then 3/$23/2<p_{0}<$ 2 and for $21<p<$ 2 we have
$\max\Big{\{}\frac{2n+p-2}{p-1}\,,\;\frac{2n-p+1}{2(2-p)}\Big{\}}={\left\{\begin{array}[]{ll}{\frac{2n+p-2}{p-1},}&{\mbox{ if }1<p\leq p_{0},}\\ {\frac{2n-p+1}{2(2-p)},}&{\mbox{ if }p_{0}\leq p<2.}\end{array}\right.}$
Theorem 4.
Let $R>\rho>0$ and let $T$ denote the torus (5). Let $p>1$ and assume that conditions (i) and (ii) of Proposition 3 are satisfied. Then
$({\rm i})$ There holds $\Delta_{p,\mathbb{H}^{n}}d\leq 0$ in the distributional sense in $T$.
$({\rm ii})$ For any $u\in C_{c}^{\infty}(T)$ there holds
$\hskip 51.21504pt\int_{T}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{T}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{p}}{d^{p}}|u|^{p}{\rm d}\xi.$
Moreover the constant in (ii) is the best possible.
Proof.
Let $p>1$. By Proposition 3 we have $\Delta_{p,\mathbb{H}^{n}}d\leq 0$ in $T\setminus S$. Hence the inequality in (ii) for any $u\in C_{c}^{\infty}(T\setminus S)$ follows from Theorem 1.
In order to extend this to any $u\in C_{c}^{\infty}(T)$ it is enough to establish that $\Delta_{p,\mathbb{H}^{n}}d\leq 0$ in the distributional sense in $T$. That is, we must prove that given a non-negative function $\phi\in{C^{\infty}_{c}}(T)$ there holds
(11) $\int_{T}|\nabla_{\mathbb{H}^{n}}d|^{p-2}\nabla_{\mathbb{H}^{n}}d\cdot\nabla_{\mathbb{H}^{n}}\phi\,{\rm d}\xi\geq 0.$
For this we shall use a standard approximation argument. Let
$q(\xi)=\sqrt{(r-R)^{2}+t^{2}}\,,\qquad\xi=(r,\omega,t)\in T,$
be the (Euclidean) distance of $\xi\in T$ to the ‘circle’ $S$. For $\epsilon>0$ small we consider a smooth function $\psi_{\epsilon}$ on $T$ such that
$\psi_{\epsilon}(\xi)={\left\{\begin{array}[]{ll}{0,}&{\mbox{ if }q(\xi)<\epsilon,}\\ {1,}&{\mbox{ if }q(\xi)>2\epsilon}\end{array}\right.}$
and $|\nabla\psi_{\epsilon}|\leq c/\epsilon$. Then $\phi_{\epsilon}:=\psi_{\epsilon}\phi$ is a non-negative smooth function in ${C^{\infty}_{c}}(T\setminus S)$ and hence, by Proposition 3,
$\int_{T}|\nabla_{\mathbb{H}^{n}}d|^{p-2}\nabla_{\mathbb{H}^{n}}d\cdot\nabla_{\mathbb{H}^{n}}\phi_{\epsilon}\,{\rm d}\xi\geq 0.$
Since $|\nabla_{\mathbb{H}^{n}}d|$ is bounded, in order to complete the proof it is enough to show that
$\int_{T}|\nabla_{\mathbb{H}^{n}}\phi_{\epsilon}-\nabla_{\mathbb{H}^{n}}\phi|\,{\rm d}\xi\longrightarrow 0\;,\quad\mbox{ as }\epsilon\to 0,$
since (11) will then follow by letting $\epsilon\to 0$.
In fact, since $|\nabla_{\mathbb{H}^{n}}u|\leq c|\nabla u|$ in $T$, it is enough to consider the Euclidean gradient. We have
$\|\nabla\phi_{\epsilon}-\nabla\phi\|_{L^{1}(T)}\leq\|(1-\psi_{\epsilon})\phi\|_{L^{1}(T)}+\|\phi\nabla\psi_{\epsilon}\|_{L^{1}(T)}.$
The first norm in the RHS tends to zero as $\epsilon\to 0$ by the Dominated Convergence Theorem. For the second one we have
$\int_{T}|\phi\nabla\psi_{\epsilon}|{\rm d}\xi$ $\leq\frac{c}{\epsilon}\int_{\{\epsilon<q(\xi)<2\epsilon\}}{\rm d}\xi$
$\leq\frac{c_{1}}{\epsilon}\epsilon^{2}$
$\to 0.$
Hence the Hardy inequality (ii) has been proved.
To establish the optimality of the constant we apply part (b) of Theorem 1. Assumption (i) is satisfied by (9). The fact that (ii) is satisfied is well known, see [BFT04, Lemma 5.2]. This completes the proof of the theorem. ∎
Remark 2.
It is evident from the argument in the above proof that if $\Omega\subset\mathbb{H}^{n}$ is a domain with $C^{2}$ boundary then the corresponding Hardy constant connot be larger than $((p-1)/p)^{p}$, provided there exists a point $\xi_{0}\in\partial\Omega$ such that $\nabla_{\mathbb{H}^{n}}d(\xi_{0})\neq 0$. This, of course, is very generic. We do not pursue this any further here, but we make two comments.
1. Assume that $\Omega$ possesses cylindrical symmetry so that $d=d(r,t)$. Suppose that at a point $\xi_{0}\in\partial\Omega$ we have $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, that is
$d_{x_{i}}+2y_{i}d_{t}=0\;,\qquad d_{y_{i}}-2x_{i}d_{t}=0\;,\qquad i=1,\ldots,n.$
Multiplying by $x_{i}$ and $y_{i}$ respectively and adding we obtain
$\sum_{i=1}^{n}(x_{i}d_{x_{i}}+y_{i}d_{y_{i}})=0,\qquad\mbox{ at the point }\xi_{0}.$
Using cylindrical coordinates $(r,\omega,t)$ we then obtain that $d_{r}(\xi_{0})=0$. By cylindrical symmetry we have
$d_{r}^{2}+d_{t}^{2}=|\nabla d|^{2}=1\;,\qquad\mbox{ in }T,$
hence $d_{t}(\xi_{0})=1$ and therefore the tangent hyperplane at $\xi_{0}$ must be parallel to the hyperplane $\{t=0\}$. Hence generically we have $\nabla_{\mathbb{H}^{n}}d\neq 0$ on $\partial\Omega$
2. Let $\Omega$ be a domain which does not necessarily possess some kind of symmetry. If $\xi_{0}\in\partial\Omega$ is a boundary point with $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, then we have
$d_{x_{i}}^{2}=4y_{i}^{2}d_{t}^{2}\;,\qquad d_{y_{i}}^{2}=4x_{i}^{2}d_{t}^{2}\;,\quad i=1,\ldots,n,\quad\mbox{ at the point }\xi_{0}.$
Since $|\nabla d|=1$, adding implies
$(1+4r^{2})d_{t}^{2}=1,\quad\mbox{ at the point }\xi_{0}.$
It follows that if in addition the tangent hyperplane at $\xi_{0}$ is parallel to the hyperplane $\{t=0\}$ (and so $d_{t}(\xi_{0})=1$), then we must necessarily have $r=0$, that is the point $\xi_{0}$ must lie on the $t$-axis.
3.2. Hardy inequalities with respect to the gauge pseudodistance
In this section we consider geometric Hardy inequalities in the Heiseberg group with respect to the gauge quasi-norm
$N(\xi)=\Big{(}(|x|^{2}+|y|^{2})^{2}+t^{2}\Big{)}^{\frac{1}{4}},\qquad\xi=(x,y,t)\in{\mathbb{H}}^{n}\,.$
For a given domain $\Omega\subset\mathbb{H}^{n}$ the induced distance to the boundary is given by
(12) $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Omega)=\inf\{N((\xi^{\prime})^{-1}\xi),\;\xi^{\prime}\in\partial\Omega\},\qquad\xi\in\Omega.$
We first consider the case where our domain is the half-space
$\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}.$
It is then easy to see that for $\xi=(x,y,t)\in\Pi_{0}$ and $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ we have
(13) $d_{N}(\xi,\xi^{\prime})=\bigg{(}\Big{(}|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}\Big{)}^{2}+\Big{(}t+2(x\cdot y^{\prime}-y\cdot x^{\prime})\Big{)}^{2}\bigg{)}^{\frac{1}{4}}.$
Lemma 5.
Let $\Pi_{0}=\{(x,y,t)\in{\mathbb{H}}^{n}:t>0\}$ and let
(14) $d_{N}(\xi)=\inf\{N((\xi^{\prime})^{-1}\xi),\;\xi^{\prime}\in\partial\Pi_{0}\}\;,\qquad\xi\in\Pi_{0},$
denote the corresponding pseudodistance to the boundary. Then, for any $\xi=(x,y,t)\in\Pi_{0}$, $d_{N}(\xi)$ depends only on $r=\sqrt{|x|^{2}+|y|^{2}}$ and $t>0$. More precisely, we have
$d_{N}(\xi)=d_{N}(r,t)={\left\{\begin{array}[]{ll}{\big{(}2r^{4}s^{2}-3tr^{2}s+t^{2}\big{)}^{\frac{1}{4}},}&{r>0,\;t>0,}\\ {t^{\frac{1}{2}},}&{r=0,\;t>0,}\end{array}\right.}$
where for fixed $r,t>0$ the real number $s\in\mathbb{R}$ is the unique solution of the equation
(15) $s^{3}+2s-\frac{t}{r^{2}}=0\;.$
Proof.
The case $r=0$ is immediate, so we assume that $r>0$. By (13) the infimum in (14) is attained at a point $(x^{\prime},y^{\prime})$ which is a critical point of the function
$F(x^{\prime},y^{\prime})=\Big{(}|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}\Big{)}^{2}+\Big{(}t+2(x\cdot y^{\prime}-y\cdot x^{\prime})\Big{)}^{2},\quad(x^{\prime},y^{\prime})\in\mathbb{R}^{2n}.$
For $i=1,\ldots,n$ we have
$F_{x_{i}^{\prime}}$ $=4\Big{(}|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}\Big{)}(x_{i}^{\prime}-x_{i})-4\Big{(}t+2(x\cdot y^{\prime}-y\cdot x^{\prime})\Big{)}y_{i}\,,$
$F_{y_{i}^{\prime}}$ $=4\Big{(}|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}\Big{)}(y_{i}^{\prime}-y_{i})+4\Big{(}t+2(x\cdot y^{\prime}-y\cdot x^{\prime})\Big{)}x_{i}\,.$
Assume now that $(x^{\prime},y^{\prime})$ is a critical point of $F$. Then necessarily $(x^{\prime},y^{\prime})\neq(x,y)$. From the last two relations we then obtain
$x_{i}(x_{i}^{\prime}-x_{i})+y_{i}(y_{i}^{\prime}-y_{i})=0,\qquad i=1,\ldots,n.$
We set
$s=\frac{t+2(x\cdot y^{\prime}-y\cdot x^{\prime})}{|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}}.$
and note that
(16) $x_{i}^{\prime}-x_{i}=sy_{i}\;,\qquad y_{i}^{\prime}-y_{i}=-sx_{i}\;,\qquad i=1,\ldots,n.$
We then have
(17) $x\cdot y^{\prime}-y\cdot x^{\prime}=\sum_{j=1}^{n}\big{[}x_{j}(y_{j}^{\prime}-y_{j})-y_{j}(x_{j}^{\prime}-x_{j})\big{]}=-r^{2}s$
and
$|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}=s^{2}\sum_{j=1}^{n}(y_{j}^{2}+x_{j}^{2})=r^{2}s^{2}.$
Hence for $i=1,\ldots,n$, we have
$F_{x_{i}^{\prime}}=4y_{i}(r^{2}s^{3}+2r^{2}s-t),\qquad F_{y_{i}^{\prime}}=-4x_{i}(r^{2}s^{3}+2r^{2}s-t)$
and we thus conclude that $s$ must solve (15).
Since the cubic equation has a unique solution, there exists a unique critical point $(x^{\prime},y^{\prime})$ of $F$ given by (16).
Finally, by (17) we have
$d_{N}^{4}(\xi)$ $=\Big{(}|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}\Big{)}^{2}+\Big{(}t+2(x\cdot y^{\prime}-x^{\prime}\cdot y)\Big{)}^{2}$
$=\Big{(}|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}\Big{)}^{2}+\Big{(}t-2r^{2}s\Big{)}^{2}$
(18) $=2r^{4}s^{2}-3tr^{2}s+t^{2},$
where we have also used (15). This completes the proof.
Proposition 6.
Let $\Pi\subset\mathbb{H}^{n}$ be an arbitrary half-space and let $d_{N}(\xi)$, $\xi\in\Pi$, denote the pseudodistance to the boundary with respect to the quasi-norm $N$. Then
$({\rm i})$ The function $d_{N}$ is $C^{1}$ in $\Pi$.
$({\rm ii})$ For any $p>1$ there holds $\Delta_{p,{\mathbb{H}}^{n}}d_{N}\leq 0$ in the distributional sense in $\Pi$.
Proof.
For simplicity we will write $d$ instead of $d_{N}$. By group action (see also [Lar16, p340]) it is enough to consider the case $\Pi=\Pi_{0}$. Also, it is preferable to work with the function
$G(r,t)=d(r,t)^{4}$
instead of $d(r,t)$. To compute the various derivatives of $G(r,t)$ we recall from Lemma 15 that
(19) $G(r,t)=2r^{4}s^{2}-3tr^{2}s+t^{2}=(t-r^{2}s)(t-2r^{2}s)\quad\mbox{ in }\Pi_{0}\setminus\{r=0\},$
where $s=s(r,t)$ is defined by (15). Since $t=r^{2}(s^{3}+2s)$ we may eliminate $t$ from (19) and we obtain
(20) $G(r,t)=r^{4}s^{4}(s^{2}+1).$
By (15) we have
(21) $s_{t}=\frac{1}{r^{2}(3s^{2}+2)}\;,\qquad s_{r}=-\frac{2s(s^{2}+2)}{r(3s^{2}+2)}.$
Hence
$G_{r}$ $=4r^{3}s^{4}(s^{2}+1)+r^{4}(6s^{5}+4s^{3})s_{r}$
$=4r^{3}s^{4}(s^{2}+1)-r^{4}(6s^{5}+4s^{3})\frac{2s(s^{2}+2)}{r(3s^{2}+2)}$
(22) $=-4r^{3}s^{4}.$
Similarly we obtain
(23) $G_{t}$ $=2r^{2}s^{3},\qquad\qquad G_{rr}=-\frac{4r^{2}s^{4}(s^{2}-10)}{3s^{2}+2}$
$G_{tt}$ $=\frac{6s^{2}}{3s^{2}+2},\qquad\quad G_{rt}=-\frac{16rs^{3}}{3s^{2}+2}.$
It is clear from (19) that to prove (i) we only need to restrict out attention near the half-line $\{r=0,\;t>0\}$.
From (15) we easily find that for fixed $t>0$ we have
$s=\frac{t^{\frac{1}{3}}}{r^{\frac{2}{3}}}-\frac{2}{3}\frac{r^{\frac{2}{3}}}{t^{\frac{1}{3}}}+o(r^{\frac{2}{3}}),\qquad\mbox{ as }r\to 0.$
It then follows from (20) that
$G(r,t)=t^{2}-3t^{\frac{4}{3}}r^{\frac{4}{3}}+o(r^{\frac{4}{3}}),\qquad\mbox{ as }r\to 0.$
For the partial derivatives of $G$ we only need to consider the first asymptotic term. From (22) and (23) we arrive at
$G_{r}(r,t)=-4t^{\frac{4}{3}}r^{\frac{1}{3}}+o(r^{\frac{1}{3}})\;,\qquad\quad G_{t}(r,t)=2t+o(1).$
It follows that
$G_{r}(0,t)=0=\lim_{r\to 0}G_{r}(r,t)\;,\qquad G_{t}(0,t)=2t=\lim_{r\to 0}G_{t}(r,t).$
Since the various $o(1)$ above are locally uniform with respect to $t>0$, we conclude that $G$ is $C^{1}$ in $\Pi_{0}$.
To prove (ii) we first note that setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (8),
(24) $\Delta_{p,\mathbb{H}^{n}}d=A^{\frac{p-4}{2}}\Big{(}A\,\Delta_{\mathbb{H}^{n}}d+\frac{p-2}{2}\big{(}d_{r}A_{r}+4r^{2}d_{t}A_{t}\big{)}\Big{)}.$
We have
$d_{r}=\frac{1}{4}G^{-\frac{3}{4}}G_{r},\qquad d_{t}=\frac{1}{4}G^{-\frac{3}{4}}G_{t},\qquad d_{rr}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{rr}$
and
$d_{tt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{t}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{tt},\qquad d_{rt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}G_{t}+\frac{1}{4}G^{-\frac{3}{4}}G_{rt}.$
Moreover
$A=d_{r}^{2}+4r^{2}d_{t}^{2}=\frac{1}{16}G^{-\frac{3}{2}}(G_{r}^{2}+4r^{2}G_{t}^{2})=G^{-\frac{3}{2}}r^{6}s^{6}(s^{2}+1),$
$A_{r}$ $=2d_{r}d_{rr}+8rd_{t}^{2}+8r^{2}d_{t}d_{rt}$
$=G^{-\frac{5}{2}}\Big{(}-\frac{3}{32}G_{r}^{3}+\frac{1}{8}G\,G_{r}G_{rr}+\frac{r}{2}G\,G_{t}^{2}-\frac{3r^{2}}{8}G_{r}G_{t}^{2}+\frac{r^{2}}{2}G\,G_{t}G_{rt}\Big{)}$
$=G^{-\frac{5}{2}}\,\frac{2r^{9}s^{12}(s^{2}+2)(s^{2}+1)}{3s^{2}+2}$
and
$A_{t}$ $=2d_{r}d_{rt}+8r^{2}d_{t}d_{tt}$
$=G^{-\frac{5}{2}}\Big{(}-\frac{3}{32}G_{r}^{2}G_{t}+\frac{1}{8}G\,G_{r}G_{rt}-\frac{3r^{2}}{8}G_{t}^{3}+\frac{r^{2}}{2}G\,G_{t}G_{tt}\Big{)}$
$=-G^{-\frac{5}{2}}\,\frac{r^{8}s^{11}(s^{2}+1)}{3s^{2}+2}.$
Combining the above we arrive at
(25) $d_{r}A_{r}+4r^{2}d_{t}A_{t}=-2\,G^{-\frac{13}{4}}\,\frac{r^{12}s^{14}(s^{2}+1)^{3}}{3s^{2}+2}.$
On the other hand in $\Pi_{0}\setminus\{r=0\}$ we have, cf. (6),
$\Delta_{{\mathbb{H}}^{n}}d$ $=d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt}$
$=\frac{1}{4}G^{-\frac{7}{4}}\Big{\{}GG_{rr}-\frac{3}{4}G_{r}^{2}+\frac{2n-1}{r}GG_{r}+4r^{2}GG_{tt}-3r^{2}G_{t}^{2}\Big{\}}$
(26) $=-G^{-\frac{7}{4}}\,\frac{r^{6}s^{8}(s^{2}+1)\big{(}(6n-2)s^{2}+4n-3\big{)}}{3s^{2}+2}.$
From (24), (25) and (26) we obtain
$\Delta_{p,\mathbb{H}^{n}}d=-G^{-\frac{3p+1}{4}}\,\frac{r^{3p}s^{3p+2}(s^{2}+1)^{\frac{p}{2}}\big{(}(6n+p-4)s^{2}+4n+p-5\big{)}}{3s^{2}+2}\leq 0,$
and the desired inequality has been proved pointwise in $\Pi_{0}\setminus\{r=0\}$ (where $d_{N}$ is smooth). To complete the proof we argue as in the proof of Theorem 4, using in particular functions $\psi_{\epsilon}$, $\epsilon>0$, as in that proof. Part (i) is also used at this point since the local boundedness of $|\nabla_{H}d|$ is required when letting $\epsilon\to 0$. ∎
Proposition 7.
Let $\Pi_{0}=\{(x,y,t)\in{\mathbb{H}}^{n}:t>0\}$ and let $d_{N}=d_{N}(r,t)$ denote the corresponding gauge pseudodistance to the boundary $\partial\Pi_{0}$ of the point $\xi=(r,\omega,t)\in\Pi_{0}$ expressed in cylindrical coordinates. Then for any fixed $r\neq 0$ we have
${\rm(i)}$ $\quad d_{N}(r,t)=\frac{t}{2r}+O(t^{3})$
${\rm(ii)}$ $\quad|\nabla_{{\mathbb{H}}^{n}}d_{N}(r,t)|=1+O(t^{2})$
as $t\to 0+$.
Proof.
For simplicity we write $d$ instead of $d_{N}$. Differentiating (18) we get
$4d^{3}d_{t}=2t-3r^{2}s+(4r^{4}s-3tr^{2})s_{t}.$
Now using the first part of (21) and the fact that $s$ solves (15) we obtain
$d_{t}=\frac{6ts^{2}-2r^{2}s+t-9r^{2}s^{3}}{4d^{3}(3s^{2}+2)}=\frac{6ts^{2}+16r^{2}s-8t}{4d^{3}(3s^{2}+2)}.$
Similarly we find that
$d_{r}=\frac{24r^{4}s^{4}+16r^{4}s^{2}-18tr^{2}s^{3}-20tr^{2}s+6t^{2}}{4d^{3}r(3s^{2}+2)}=\frac{40r^{2}ts-32r^{4}s^{2}-12t^{2}}{4d^{3}r(3s^{2}+2)}.$
We now let $t\to 0+$. From (15) we find
$s=\frac{t}{2r^{2}}-\frac{t^{3}}{16r^{6}}+O(t^{5}).$
Plugging this in (18) we have
$d^{4}(r,t)=\frac{t^{4}}{16r^{4}}-\frac{t^{6}}{64r^{8}}+O(t^{8})$
and (i) follows. We then also have
$\frac{1}{d^{6}(r,t)}=\frac{64r^{6}}{t^{6}}+O(\frac{1}{t^{4}})$
and combining the above we obtain
$d_{t}^{2}=\frac{1}{4r^{2}}+O(t^{2})\;,\qquad d_{r}^{2}=\frac{t^{2}}{4r^{4}}+O(t^{4}).$
We thus conclude that
$|\nabla_{{\mathbb{H}}^{n}}d|^{2}=d_{r}^{2}+4r^{2}d_{t}^{2}=1+O(t^{2}),$
as required. ∎
Theorem 8.
Let $p>1$ and $\Pi$ be an arbitrary half-space in $\mathbb{H}^{n}$. Let $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Pi)$ denote the corresponding pseudodistance of $\xi\in\Pi$ to the boundary $\partial\Pi$. Then there holds
$\int_{\Pi}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Pi}\frac{|\nabla_{{\mathbb{H}}^{n}}d_{N}|^{p}}{d_{N}^{p}}|u|^{p}{\rm d}\xi\;,\quad u\in{C^{\infty}_{c}}(\Pi).$
Moreover the constant is the best possible.
Proof.
Action by an appropriate group element reduces the proof to the case $\Pi=\Pi_{0}=\{(x,y,t):t>0\}$. The validity of the Hardy inequality is a consequence of Theorem 1 (a) and Proposition 6. The sharpness of the constant follows from the second part of Theorem 1 (b) and Proposition 7. ∎
In case $p=2$ we will extend the above to the case of a bounded convex polytope. For this we will need the following lemma where, as above, $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:\;t>0\}$.
Lemma 9.
Any point $\xi\in\Pi_{0}$ has a unique nearest boundary point $(x^{\prime},y^{\prime},0)\in\partial\Pi$. Moreover, given a point $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ and $\rho>0$, there exists a unique point $\xi=(x,y,t)\in\Pi_{0}$ whose nearest boundary point is $\xi^{\prime}$ and for which $d_{N}(\xi)=\rho$.
Proof.
We have already seen in the proof of Lemma 15 that given $\xi\in\Pi_{0}$ the nearest boundary point $\xi^{\prime}\in\partial\Pi$ is uniquely defined.
Suppose now that a point $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ and $\rho>0$ are given. Assume that $\xi\in\Pi_{0}$ has $\xi^{\prime}$ as its nearest boundary point and that $d_{N}(\xi)=\rho$. Denoting $r^{2}=|x|^{2}+|y|^{2}$ and $r^{\prime 2}=|x^{\prime}|^{2}+|y^{\prime}|^{2}$ we have from (16) that
$r^{\prime 2}=(1+s^{2})r^{2}\,,$
where $s>0$ is defined in terms of $r,t$ by (15). We also have (cf. (20)) $\rho^{4}=r^{4}s^{4}(1+s^{2})$. We thus conclude that
$\frac{r^{\prime 4}}{\rho^{4}}=\frac{1+s^{2}}{s^{4}},$
and this relation uniquely determines $s>0$. Now going back to (16) we obtain
$x_{i}=\frac{x_{i}^{\prime}-sy_{i}^{\prime}}{1+s^{2}},\qquad y_{i}=\frac{sx_{i}^{\prime}+y_{i}^{\prime}}{1+s^{2}}.$
We also have $t=r^{2}(s^{3}+2s)$, hence the point $\xi=(x,y,t)\in\Pi_{0}$ has been uniquely determined. It is not difficult now to see that this point has indeed $(x^{\prime},y^{\prime},0)$ as its nearest boundary point and $d_{N}(\xi)=\rho$. This completes the proof. ∎
Remark 3.
Let us point out that the convexity of a set $\Omega\subset G$ in a stratified group $G$ is a genuine geometric notion in the sense that it is invariant under left translations; i.e., if $\Omega\subset G$ is convex and $g\in G$, then $g\Omega$ is also convex.
Theorem 10.
Let $\Omega\subset\mathbb{H}^{n}$ be a bounded convex polytope and let $d_{N}(\xi)$, $\xi\in\Omega$, denote the corresponding gauge pseudodistance to the boundary. Then
${\rm(i)}$ $\quad\Delta_{{\mathbb{H}}^{n}}d_{N}\leq 0\;\mbox{ in the distributional sense in }\Omega.$
${\rm(ii)}$ The Hardy inequality
$\hskip 42.67912pt\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}\xi\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{{\mathbb{H}}^{n}}d_{N}|^{2}}{d_{N}^{2}}u^{2}{\rm d}\xi\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
is valid.
Proof.
Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
$A_{k}=\{\xi\in\Omega:d(\xi)={\rm dist}_{N}(\xi,E_{k}),\},\qquad k=1,\ldots,m.$
Hence the sets $A_{k}$ have pairwise disjoint interiors and $\cup_{k=1}^{m}A_{k}=\Omega$. Let $\Pi_{k}$, $k=1,\ldots,m$, denote the half-spaces determined by $\Omega$ so that
$E_{k}\subset\partial\Pi_{k},\;\;k=1,\ldots,m,\quad\mbox{ and}\qquad\Omega=\bigcap_{k=1}^{m}\Pi_{k}.$
We claim that
(27) $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
Indeed, suppose to the contary that (27) is not true for some $k$ and some $\xi_{0}\in A_{k}$. Then there exists $\xi_{0}^{\prime}\in\partial\Pi_{k}\setminus E_{k}$ such that
${\rm dist}_{N}(\xi_{0},\partial\Pi_{k})=d_{N}(\xi_{0},\xi_{0}^{\prime}).$
By Lemma 9 there exists a continuous curve joining $\xi_{0}$ and $\xi_{0}^{\prime}$ with the property that each point on the curve has $\xi_{0}^{\prime}$ as its nearest point on $\partial\Pi_{k}$. Since $\xi_{0}^{\prime}\not\in E_{k}$, that curve necessarily intersects some $A_{j}$, $j\neq k$, which is a contradiction.
It immediately follows from (27) that
$d_{N}(\xi)=\min_{1\leq k\leq m}d_{k}(\xi),\qquad\xi\in\Omega.$
To prove (i) we fix a non-negative function $\phi\in{C^{\infty}_{c}}(\Omega)$ and we aim to show that
$\int_{\Omega}\nabla_{{\mathbb{H}}^{n}}d\cdot\nabla_{{\mathbb{H}}^{n}}\phi\,{\rm d}\xi\geq 0.$
We recall the divergence theorem in the stratified setting: if the vector field $F$ takes values in the first stratum and sufficient regularity is assumed then
[…]
where $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ and $\nu=(\nu_{x},\nu_{y},\nu_{t})$ denotes the usual outer normal.
Let $k\in\{1,\ldots,m\}$ be fixed. Integrating by parts and using Proposition 6 we obtain
$\int_{A_{k}}\nabla_{{\mathbb{H}}^{n}}d\cdot\phi\,{\rm d}\xi=$ $-\int_{A_{k}}\phi\,\Delta_{{\mathbb{H}}^{n}}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi\,\nabla_{{\mathbb{H}}^{n}}d\cdot\nu_{H}dS$
$\geq$ $\int_{\partial A_{k}}\!\!\phi\,\nabla_{{\mathbb{H}}^{n}}d_{k}\cdot\nu_{H}dS.$
We then add over all $k=1,\ldots,m$.
Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$. Let us fix such a set $A_{j}$, $j\neq k$. Denoting by $\nu_{H}$ the horizontal normal vector which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S$ from $A_{k}$ and $A_{j}$ add up to
[…]
The surface $S$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$. We therefore have
$(\nabla_{{\mathbb{H}}^{n}}\,d_{k}-\nabla_{{\mathbb{H}}^{n}}\,d_{j})\cdot\nu_{H}$
$=\big{(}\nabla_{\!x}\,d_{k}-\nabla_{\!x}\,d_{j}+2d_{k,t}y-2d_{j,t}y,\nabla_{\!y}\,d_{k}-\nabla_{\!y}\,d_{j}-2d_{k,t}x+2d_{j,t}x\big{)}$
$\quad\quad\cdot\big{(}\nu_{x}+2y\nu_{t},\nu_{y}-2x\nu_{t}\big{)}$
$=\big{(}\lambda\nu_{x}+2\lambda\nu_{t}y,\lambda\nu_{y}-2\lambda\nu_{t}x\big{)}\cdot\big{(}\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x\big{)}$
$=\lambda\Big{(}|\nu_{x}+2\nu_{t}y|^{2}+|\nu_{y}-2\nu_{t}x|^{2}\Big{)}$
$\geq 0.$
Combining the above completes the proof of (i). Part (ii) is an immediate consequence of part (i) and Theorem 1. ∎
As an immediate consequence of Theorem 10 we obtain the geometric uncertainty principle on the convex set $\Omega\subset{\mathbb{H}}^{n}$ with respect to the gauge pseudodistance on ${\mathbb{H}}^{n}$.
Corollary 11.
Let $D\subset\mathbb{H}^{n}$ be either a bounded convex polytope or an arbitary half-space in ${\mathbb{H}}^{n}$ and let
$d_{N}(\xi)={\rm dist}_{N}(\xi,\partial D)$
denote the corresponding pseudodistance of $\xi\in D$ to the boundary $\partial D$. Then for any $u\in C_{c}^{\infty}(D)$ we have
$\left(\int_{D}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}\xi\right)^{\frac{1}{2}}\left(\int_{D}d_{N}^{2}u^{2}{\rm d}\xi\right)^{\frac{1}{2}}\geq\frac{1}{2}\int_{D}u^{2}{\rm d}\xi\,.$
Proof.
A combination of Theorem 10, Part (ii) and of the Cauchy-Schwarz inequality yields
$\left(\int_{D}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}\xi\right)\left(\int_{D}d_{N}^{2}u^{2}{\rm d}\xi\right)$$\geq$ $\frac{1}{4}\left(\int_{D}\frac{u^{2}}{d_{N}^{2}}{\rm d}\xi\right)\left(\int_{D}d_{N}^{2}u^{2}{\rm d}\xi\right)$
$\geq$ $\frac{1}{4}\left(\int_{D}u^{2}{\rm d}\xi\right)^{2}.$
4. Hardy inequalities on stratified groups of step two
In this section we consider stratified groups of step two. If $G\equiv\mathbb{R}^{n}$ is such a group with (cf. (2)) $\text{dim}(V_{1})=m<n$, then each element $g\in G$ can be written as
$g=(g^{(1)},g^{(2)})=(g_{1},\cdots,g_{m},g_{m+1},\cdots,g_{n}),$
where $g^{(1)}\in\mathbb{R}^{m}$ and $g^{(2)}\in\mathbb{R}^{n-m}$ belong in the first and the second stratum of $G$, respectively. It is known that the group law has the form
(28) $(g^{\prime}g)_{i}=\left\{\begin{array}[]{ll}g_{i}+g^{\prime}_{i}\,,&i=1,\cdots,m,\\ g_{i}+g^{\prime}_{i}+\frac{1}{2}\langle B^{(i)}g^{\prime(1)},g^{(1)}\rangle\,,&i=m+1,\cdots,n,\end{array}\right.\,$
[…]
(29) $g^{\prime}g=(g^{(1)}+g^{\prime(1)},g^{(2)}+g^{\prime(2)}+\frac{1}{2}\langle Bg^{\prime(1)},g^{(1)}\rangle)\,,$
where $\langle Bg^{(1)},g^{\prime(1)}\rangle$ denotes the $(n-m)$-tuple
$(\langle B^{(m+1)}g^{(1)},g^{\prime(1)}\rangle,\cdots,\langle B^{(n)}g^{(1)},g^{\prime(1)}\rangle)\,.$
The inverse element is then given by
$(g^{(1)},g^{(2)})^{-1}=(-g^{(1)},-g^{(2)}+\frac{1}{2}\langle Bg^{(1)},g^{(1)}\rangle)\,.$
We note that the (anisotropic) dilations on a stratified group $G$ of step two are given by the maps $\delta_{\lambda}$, $\lambda>0$, defined by
$\delta_{\lambda}((g^{(1)},g^{(2)}))=(\lambda g^{(1)},\lambda^{2}g^{(2)}).$
In the first part of this section we prove some results on the concavity, in the sense of [DGN03, LMS03], of the Euclidean distance to the boundary on a convex set $\Omega\subset G$. This, combined with results for [DGN03], yields Theorem C of the introduction and in particular the $L^{2}$ Hardy inequality on convex domains $\Omega\subset G$.
4.1. On the distance function from the boundary of bounded convex domains in stratified groups
To develop the subsequent analysis we first need to clarify the notions of convexity of sets and functions in the stratified setting. Even though, as mentioned above, these notions were introduced at the same time in [DGN03] and in [LMS03], here we adopt the notation of [DGN03] since in [LMS03] these notions are developed in the viscosity sense, while for us the weak sense is more suitable. To this end let us first introduce the following auxiliary notion.
Let $G$ be a stratified group with $\text{dim}(G)=n$ and $\text{dim}(V_{1})=m$. Given a point $g\in G$ the horizontal plane $H_{g}$ passing through $g$ is defined by
$H_{g}=L_{g}(\exp(V_{1}\oplus\{0\}))\,,$
where $L_{g}$ denotes the left translation by $g\in G$ and $\exp:\mathfrak{g}\rightarrow G$ is the exponential map for the group $G$. In particular, we have that
$H_{e}=\exp(V_{1}\oplus\{0\})\,,$
where $e\in G$ is the identity element of $G$.
Following [DGN03], for given $g,g^{\prime}\in G$ and $\lambda\in[0,1]$ we denote by $g_{\lambda}$ the anisotropic analogue of the standard Euclidean convex combination, that is
$g_{\lambda}=g_{\lambda}(g;g^{\prime}):=g\delta_{\lambda}(g^{-1}g^{\prime}).$
The following definition was given in [DGN03, Definition 5.5].
Definition 12.
A function $u:G\rightarrow(-\infty,\infty]$ is called weakly $H$-convex if $\{g\in G:u(g)=\infty\}\neq G$, and if for every $g\in G$ and $g^{\prime}\in H_{g}$ one has
$u(g_{\lambda})\leq u(g)+\lambda(u(g^{\prime})-u(g))\,,\qquad\lambda\in[0,1].$
The notion of a weakly $H$-concave function can be defined accordingly.
In [DGN03, Definition 7.1] the authors introduced the following definition of convexity of sets in the stratified setting.
Definition 13.
A subset $\Omega\subset G$ of a stratified group $G$ is called weakly $H$-convex if for any $g\in\Omega$ and for any $g^{\prime}\in\Omega\cap H_{g}$ one has $g_{\lambda}\in\Omega$ for every $\lambda\in[0,1].$
Remark 4.
It is easy to prove that if $\Omega\subset\mathbb{R}^{n}$ is convex in the Euclidean sense then $\Omega$ is a weakly $H$-convex set in a stratified group $G\equiv\mathbb{R}^{n}$ of step two. To see this we first observe that by the identification
$\exp(g_{1}X_{1}+\cdots+g_{n}X_{n})=(g_{1},\cdots,g_{n})\,,$
between $G$ and the corresponding Lie algebra $\mathfrak{g}$ via the exponential map, we have $g\in H_{e}$ if and only if $g$ is of the form $g=(g_{1},\cdots,g_{m},0,\cdots,0)$. Since $H_{g}=L_{g}H_{e}$, we obtain from (29) that $g^{\prime}\in H_{g}$ if and only if $g^{\prime}$ is of the form
(30) $g^{\prime}=(g^{\prime(1)},g^{\prime(2)})=(g^{(1)}+v^{(1)},g^{(2)}+\frac{1}{2}\langle Bg^{(1)},v^{(1)}\rangle)$
for some $v^{(1)}\in\mathbb{R}^{m}$. Suppose now that $g\in\Omega$ and let $g^{\prime}\in\Omega\cap H_{g}$. Then $g^{-1}g^{\prime}\in H_{e}$, which in turn implies that
$\delta_{\lambda}(g^{-1}g^{\prime})=(\lambda(g^{\prime(1)}-g^{(1)}),0)\,.$
So
$g_{\lambda}$ $=$ $(g^{(1)}+\lambda(g^{\prime(1)}-g^{(1)}),g^{(2)}+\frac{1}{2}\langle Bg^{(1)},\lambda(g^{\prime(1)}-g^{(1)}))$
$=$ $(g^{(1)}+\lambda(g^{\prime(1)}-g^{(1)}),g^{(2)}+\frac{1}{2}\lambda\langle Bg^{(1)},v^{(1)}\rangle\,)$
since by (30) we have $g^{\prime(1)}=g^{(1)}+v^{(1)}$, for some $v\in H_{e}$. Using (30) we conclude that
(31) $g_{\lambda}=(1-\lambda)g+\lambda g^{\prime}\,.$
Hence if $\Omega$ is convex (in the Euclidean sense) it is also weakly $H$-convex.
In the following example we show that the distance to a hyperplane with respect to the quasi-norm (1) is not weakly $H$-concave.
Example 1.
Let $\mathbb{H}^{n}$ be the Heisenberg group and $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}$. We shall prove that the the distance $d_{N}$ (cf. (12)) is not necessarily weakly $H$-concave.
Actually, we shall show that the weak $H$-concavity fails in a neighbourhood of any boundary point. Indeed, let $\xi=(x,y,t)\in\Pi_{0}$, and for fixed $\alpha>0$ let $\xi^{\prime}=(x^{\prime},y^{\prime},t)=(ax,ay,t).$ Then $\xi^{-1}\xi^{\prime}=(x^{\prime}-x,y^{\prime}-y,0)\in H_{e}$, hence $\xi^{\prime}\in H_{\xi}$. Given $\lambda\in(0,1)$ we have by (31)
$\xi_{\lambda}=\big{(}\lambda x^{\prime}+(1-\lambda)x,\;\lambda y^{\prime}+(1-\lambda)y,\;t\big{)}.$
We use cylindrical coordinates (cf. Section 3) and write
$\xi=(r,\omega,t),\quad\xi^{\prime}=(r^{\prime},\omega,t),\quad\xi_{\lambda}=(r_{\lambda},\omega,t)$
Assume now for contradiction that $d_{N}$ is weakly $H$-convex. Then
$d_{N}(r_{\lambda},t)\geq(1-\lambda)d_{N}(r,t)+\lambda d_{N}(r^{\prime},t).$
Using the asymptotics of Proposition 7 we then have
$\frac{t}{2r_{\lambda}}\geq(1-\lambda)\frac{t}{2r}+\lambda\frac{t}{2r^{\prime}}+O(t^{3}),\qquad\mbox{ as }t\to 0+.$
Hence
$\frac{1}{r_{\lambda}}\geq(1-\lambda)\frac{1}{r}+\lambda\frac{1}{r^{\prime}}\;,$
which contradicts the strict convexity of the function $1/r$. We note that the above argument can be implemented in a small neighbourood of any boundary point $\xi_{0}\in\partial\Pi_{0}$.
4.2. Hardy inequalities with respect to the Euclidean distance
In this section we prove that the Euclidean distance to the boundary on a convex, bounded domain $\Omega$ is weakly $H$-concave and superharmonic. This provides an alternative proof of the $L^{2}$-Hardy inequality for such domains.
Theorem 14.
Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a convex, in the Euclidean sense, bounded domain in $G$. Then the Euclidean distance to the boundary is a weakly $H$-concave function on $\Omega$.
Proof.
Let $\Omega$ be as in the hypothesis and let $g,g^{\prime}\in\Omega$, with $g^{\prime}\in H_{g}$. We want to show that for any $\lambda\in[0,1]$ we have
(32) $d(g\delta_{\lambda}(g^{-1}g^{\prime}))\geq(1-\lambda)d(g)+\lambda d(g^{\prime})\,.$
Notice that showing $B_{r_{\lambda}}(g_{\lambda})\subset\Omega$, where $B_{r_{\lambda}}(g_{\lambda})$ is the Euclidean ball of radius $r_{\lambda}=(1-\lambda)d(g)+\lambda d(g^{\prime})$ centered at $g_{\lambda}$, we would have the desired inequality (32). Let $h\in B_{r_{\lambda}}(g_{\lambda})$. Then $|h-g_{\lambda}|=\rho\leq r_{\lambda}$. We define
$v=\frac{h-g_{\lambda}}{\rho}\,,\qquad g_{1}=g+\rho_{1}v\,,\qquad g^{\prime}_{1}=g^{\prime}+\rho_{2}v,$
where
$\rho_{1}:=\frac{d(g)}{(1-\lambda)d(g)+\lambda d(g^{\prime})}\rho\,,\quad\text{and}\quad\rho_{2}:=\frac{d(g^{\prime})}{(1-\lambda)d(g)+\lambda d(g^{\prime})}\rho\,.$
Then $g_{1}\in B_{d(g)}(g)\subset\Omega$ and $g^{\prime}_{1}\in B_{d(g^{\prime})}(g^{\prime})\subset\Omega$, since $|v|=1$, $\rho_{1}\leq d(g)$, and $\rho_{2}\leq d(g^{\prime})$. Recalling also (31) we then have
$(1-\lambda)g_{1}+\lambda g^{\prime}_{1}$ $=$ $(1-\lambda)(g+\rho_{1}v)+\lambda(g^{\prime}+\rho_{2}v)$
$=$ $(1-\lambda)g+\lambda g^{\prime}+(1-\lambda)\rho_{1}v+\lambda\rho_{2}v$
$=$ $g_{\lambda}+\rho v$
$=$ $h\,,$
where the last inequality follows by the choice of $v$. Hence, by the Euclidean convexity of $\Omega$, we have $h\in\Omega$, and the proof is complete. ∎
From Theorem 14 we immediately obtain the following result; the sharpness of the constant 1/4 follows under the hypotheses of part (b) of Theorem 1.
Theorem 15.
Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
${\rm(i)}$ $\quad\Delta_{H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
${\rm(ii)}$ The Hardy inequality
$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{2}{\rm d}g\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{H}d|^{2}}{d^{2}}u^{2}{\rm d}g\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
is valid.
Proof.
By Theorem 14 the distance function is weakly $H$-concave. Let $X_{1},\ldots,$ $X_{m}$ be the vector fields that generate the first stratum $V_{1}$ of the corresponding Lie algebra. By Theorem [DGN03, Theorem 8.1] each $X_{k}^{2}d$, $k=1,\ldots,m$, is a non-positive Radon measure on $\Omega$; this proves (i). Part (ii) now follows from Theorem 1. ∎

### v2

@@
 We prove $L^{p}$-Hardy inequalities with distance to the boundary for domains in the Heisenberg group ${\mathbb{H}}^{n}$, $n\geq 1$.
-Our results are based on a geometric condition.
+Our results are based on a certain geometric condition.
 This is first implemented for the Euclidean distance in certain non-convex domains.
-It is also implemented on half-spaces and convex polytopes for the distance defined by the gauge quasi-norm on ${\mathbb{H}}^{n}$ related to the fundamental solution of the horizontal Laplacian.
+It is then implemented for the distance defined by the gauge quasi-norm related to the fundamental solution of the horizontal Laplacian when the domain is a half-space or a convex polytope.
+Finally it is implemented for the Carnot-Carathéodory distance on half-spaces and arbitrary bounded convex domains of ${\mathbb{H}}^{n}$.
+In all cases the constant $((p-1)/p)^{p}$ is obtained.
 In the more general context of a stratified Lie group of step two we study the superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an alternative proof for the $L^{2}$-Hardy inequality on convex domains.
-In all cases the constants are shown to be sharp.
 Key words and phrases:
-Geometric $L^{p}$-Hardy inequalities; gauge pseudodistance; Heisenberg group; stratified groups of step two; horizontal convexity
+Geometric $L^{p}$-Hardy inequalities; gauge pseudodistance; Carnot-Carathéodory distance; Heisenberg group; stratified groups of step two; horizontal convexity
 2020 Mathematics Subject Classification:
@@
 where $\nabla_{H}$ denotes the horizontal gradient on $H$. For other results in this direction see also [Rus18].
-Our main interest in this work is to prove subelliptic geometric Hardy inequalities with best constant in the Heisenberg group ${\mathbb{H}}^{n}$ but also on any stratified group of step two.
+Our main interest in this work is to prove subelliptic geometric Hardy inequalities with best constant on the Heisenberg group ${\mathbb{H}}^{n}$, but also on any stratified group of step two.
 Our approach is based on the general method of [BFT04] and in particular in the $L^{p}$-superharmonicity of the distance function.
@@
 $\int_{T}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{T}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{p}}{d^{p}}|u|^{p}{\rm d}\xi\;,\qquad u\in{C^{\infty}_{c}}(T).$
-Another class of distances in a stratified group consists of those induced by any homogenous quasi-norm. In the case of the Heisenberg group $\mathbb{H}^{n}$ one important such quasi-norm is
+Another class of distances in a stratified group consists of those induced by a homogenous quasi-norm.
+In the case of the Heisenberg group $\mathbb{H}^{n}$ there are two particularly important such (quasi)-norms.
+The first is
 (1) $N(\xi)=\Big{(}(|x|^{2}+|y|^{2})^{2}+t^{2}\Big{)}^{\frac{1}{4}},\qquad\xi=(x,y,t)\in{\mathbb{H}}^{n}.$
-We recall if we set $Q=2n+2$,the homogeneous dimension of $\mathbb{H}^{n}$, then up to a multiplicative constant, $N^{2-Q}$ is the fundamental solution of the sub-Laplacian on the Heisenberg group ${\mathbb{H}}^{n}$ [Fol73].
-Denoting by $d_{N}$ the distance to the boundary incuded by $N$, cf.
-(12) below, we have the following
+We note that $N^{2-Q}$, with $Q=2n+2$ being the homogeneous dimension of $\mathbb{H}^{n}$, is (up to a multiplicative constant) the fundamental solution of the sub-Laplacian on ${\mathbb{H}}^{n}$ [Fol73].
+A second important homogeneous quasi-norm, which respects the sub-Riemannian geometry of ${\mathbb{H}}^{n}$, is the one arising from the Carnot-Carathéodory distance, denoted here by $\rho$.
+Let us recall the definition of the Carnot-Carathéodory distance between two points $x,y\in\mathbb{R}^{n}$: For a family $\{X_{1},\cdots,X_{m}\}$ of vector fields, the Carnot-Carathéodory distance between $x,y$ is given by
+(2) $\rho(x,y)=\inf_{\gamma\in\mathcal{C}_{x,y}}\{{\rm length}(\gamma)\}\,,$
+where $\mathcal{C}_{x,y}$ is the set of horizontal curves joining $x$ to $y$.
+In the case of a stratified group when the set $\{X_{1},\cdots,X_{m}\}$ coincides with the first stratum $V_{1}$ of its Lie algebra (see Section 4 for related definitions) we always have $d(x,y)<\infty$, and the Carnot-Carathéodory distances induced by different choices of $V_{1}$ are equivalent [Pa89, Section 1.3].
+Denoting by $d_{N}$ the distance to the boundary induced by $N$, cf. (13) below, we have the following
 Theorem B. (i) Let $p>1$ and let $D\subset{\mathbb{H}}^{n}$ be a half-space. There holds
@@
 (ii) In case $p=2$ the above inequality is also valid for any bounded convex polytope.
-The constant is the best possible in both cases.
+Moreover the constant is the best possible in both cases.
+Denoting by $d_{\rho}$ the Carnot-Carathéodory distance to the boundary, we obtain the following
+Theorem C. Let $p>1$ and let $D\subset{\mathbb{H}}^{n}$ be a bounded and convex domain or a half-space in ${\mathbb{H}}^{n}$. There holds
+$\int_{D}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\left(\frac{p-1}{p}\right)^{p}\int_{D}\frac{|u|^{p}}{d_{\rho}^{p}}\,{\rm d}\xi\;,\qquad u\in{C^{\infty}_{c}}(D)\,.$
+In the case of the half-space the constant is sharp.
+The appereance of the factor $|\nabla_{{\mathbb{H}}^{n}}d_{N}|$ in Theorem B – while a similar factor does not appear in Theorem C – should be seen in the light of a result of [MSC01] a special case of which states that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e.
+A more interesting difference between the above two homogeneous norms relates to nearest boundary points in the case of a subspace $\Pi\subset{\mathbb{H}}^{n}$.
+The correspondence
+$\Pi\ni\xi\longleftrightarrow(\xi^{\prime},s)\in\partial\Pi\times\mathbb{R}_{+}$
+where $s={\rm dist}(\xi,\partial\Pi)=d(\xi,\xi^{\prime})$ is different in the two cases.
+In the case of the quasi-norm (1) the above is a simple 1-1 correspondance, as in the Euclidean case, cf.
+Proposition 9.
+However in the case of the Carnot-Carathéodory distance the situation is entirely different; see Proposition 16.
 In Section 4 we extend our setting to that of an arbitrary stratified group of step two.
-In this setting a certain notion of convexity plays a central role.
+In this general setting, a certain notion of convexity plays a central role.
 There are various notions of convexity in the sub-Riemannian setting and their properties can vary significantly [DGN03, DLZ24, LMS03].
@@
 Exploring properties of the so-called weakly $H$-concave functions (see Definition 20) and using a result from [DGN03] we prove the following theorem, part (iii) of which is contained in [RSS20].
-Theorem C. Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
+Theorem D. Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
 […]
@@
 Proposition 6.
-Let $\Pi\subset\mathbb{H}^{n}$ be an arbitrary half-space and let $d_{N}(\xi)$, $\xi\in\Pi$, denote the pseudodistance to the boundary with respect to the quasi-norm $N$. Then
-$({\rm i})$ The function $d_{N}$ is $C^{1}$ in $\Pi$.
-$({\rm ii})$ For any $p>1$ there holds $\Delta_{p,{\mathbb{H}}^{n}}d_{N}\leq 0$ in the distributional sense in $\Pi$.
+Let $\Pi\subset\mathbb{H}^{n}$ be an arbitrary half-space and let $d_{N}(\xi)$, $\xi\in\Pi$, denote the pseudodistance to the boundary with respect to the quasi-norm $N$.
+Then for any $p>1$ there holds $\Delta_{p,{\mathbb{H}}^{n}}d_{N}\leq 0$ in the distributional sense in $\Pi$.
 Proof.
@@
 $G_{tt}$ $=\frac{6s^{2}}{3s^{2}+2},\qquad\quad G_{rt}=-\frac{16rs^{3}}{3s^{2}+2}.$
-It is clear from (19) that to prove (i) we only need to restrict out attention near the half-line $\{r=0,\;t>0\}$.
-From (15) we easily find that for fixed $t>0$ we have
-$s=\frac{t^{\frac{1}{3}}}{r^{\frac{2}{3}}}-\frac{2}{3}\frac{r^{\frac{2}{3}}}{t^{\frac{1}{3}}}+o(r^{\frac{2}{3}}),\qquad\mbox{ as }r\to 0.$
-It then follows from (20) that
-$G(r,t)=t^{2}-3t^{\frac{4}{3}}r^{\frac{4}{3}}+o(r^{\frac{4}{3}}),\qquad\mbox{ as }r\to 0.$
-For the partial derivatives of $G$ we only need to consider the first asymptotic term. From (22) and (23) we arrive at
-$G_{r}(r,t)=-4t^{\frac{4}{3}}r^{\frac{1}{3}}+o(r^{\frac{1}{3}})\;,\qquad\quad G_{t}(r,t)=2t+o(1).$
-It follows that
-$G_{r}(0,t)=0=\lim_{r\to 0}G_{r}(r,t)\;,\qquad G_{t}(0,t)=2t=\lim_{r\to 0}G_{t}(r,t).$
-Since the various $o(1)$ above are locally uniform with respect to $t>0$, we conclude that $G$ is $C^{1}$ in $\Pi_{0}$.
-To prove (ii) we first note that setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (8),
+Setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (9),
 (25) $\Delta_{p,\mathbb{H}^{n}}d=A^{\frac{p-4}{2}}\Big{(}A\,\Delta_{\mathbb{H}^{n}}d+\frac{p-2}{2}\big{(}d_{r}A_{r}+4r^{2}d_{t}A_{t}\big{)}\Big{)}.$
@@
 In case $p=2$ we will extend the above to the case of a bounded convex polytope. For this we will need the following lemma where, as above, $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:\;t>0\}$.
-Lemma 9.
+Proposition 9.
 Any point $\xi\in\Pi_{0}$ has a unique nearest boundary point $(x^{\prime},y^{\prime},0)\in\partial\Pi$.
@@
 $\hskip 42.67912pt\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}\xi\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{{\mathbb{H}}^{n}}d_{N}|^{2}}{d_{N}^{2}}u^{2}{\rm d}\xi\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
-is valid.
+is valid, and the constant $1/4$ is sharp.
 Proof.
-Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
-$A_{k}=\{\xi\in\Omega:d(\xi)={\rm dist}_{N}(\xi,E_{k}),\},\qquad k=1,\ldots,m.$
+For simplicity we write $d$ instead of $d_{N}$. Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
+$A_{k}=\{\xi\in\Omega:d(\xi)={\rm dist}(\xi,E_{k}),\},\qquad k=1,\ldots,m.$
 Hence the sets $A_{k}$ have pairwise disjoint interiors and $\cup_{k=1}^{m}A_{k}=\Omega$. Let $\Pi_{k}$, $k=1,\ldots,m$, denote the half-spaces determined by $\Omega$ so that
 $E_{k}\subset\partial\Pi_{k},\;\;k=1,\ldots,m,\quad\mbox{ and}\qquad\Omega=\bigcap_{k=1}^{m}\Pi_{k}.$
-We claim that
-(27) $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
-Indeed, suppose to the contary that (27) is not true for some $k$ and some $\xi_{0}\in A_{k}$. Then there exists $\xi_{0}^{\prime}\in\partial\Pi_{k}\setminus E_{k}$ such that
-${\rm dist}_{N}(\xi_{0},\partial\Pi_{k})=d_{N}(\xi_{0},\xi_{0}^{\prime}).$
-By Lemma 9 there exists a continuous curve joining $\xi_{0}$ and $\xi_{0}^{\prime}$ with the property that each point on the curve has $\xi_{0}^{\prime}$ as its nearest point on $\partial\Pi_{k}$.
-Since $\xi_{0}^{\prime}\not\in E_{k}$, that curve necessarily intersects some $A_{j}$, $j\neq k$, which is a contradiction.
+We then have
+(28) $d(\xi)={\rm dist}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
 It immediately follows from (28) that
-$d_{N}(\xi)=\min_{1\leq k\leq m}d_{k}(\xi),\qquad\xi\in\Omega.$
-To prove (i) we fix a non-negative function $\phi\in{C^{\infty}_{c}}(\Omega)$ and we aim to show that
-$\int_{\Omega}\nabla_{{\mathbb{H}}^{n}}d\cdot\nabla_{{\mathbb{H}}^{n}}\phi\,{\rm d}\xi\geq 0.$
+$d(\xi)=\min_{1\leq k\leq m}d_{k}(\xi),\qquad\xi\in\Omega.$
+We fix a non-negative function $\phi\in{C^{\infty}_{c}}(\Omega)$ and we aim to show that
+(29) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi\,{\rm d}\xi=\sum_{k=1}^{m}\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi\,{\rm d}\xi\geq 0.$
 We recall the divergence theorem in the stratified setting: if the vector field $F$ takes values in the first stratum and sufficient regularity is assumed then
 […]
-where $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ and $\nu=(\nu_{x},\nu_{y},\nu_{t})$ denotes the usual outer normal.
-Let $k\in\{1,\ldots,m\}$ be fixed. Integrating by parts and using Proposition 6 we obtain
-$\int_{A_{k}}\nabla_{{\mathbb{H}}^{n}}d\cdot\phi\,{\rm d}\xi=$ $-\int_{A_{k}}\phi\,\Delta_{{\mathbb{H}}^{n}}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi\,\nabla_{{\mathbb{H}}^{n}}d\cdot\nu_{H}dS$
-$\geq$ $\int_{\partial A_{k}}\!\!\phi\,\nabla_{{\mathbb{H}}^{n}}d_{k}\cdot\nu_{H}dS.$
-We then add over all $k=1,\ldots,m$.
+where
+(30) $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$
+and $\nu=(\nu_{x},\nu_{y},\nu_{t})$ denotes the usual outer normal vector.
+We cannot directly apply integration by part to each of the integrals in the RHS of (29) due to the fact that, as is seen from Lemma 5, there is a halfline $L_{k}\subset\Pi_{k}$ on which each function $d_{k}$ is not differentiable (the halfline $L_{k}$ is the image of the halfline $\{(0,0,t):t>0\}$ under the group action that maps $\Pi_{0}$ onto $\Pi_{k}$).
+We define the set
+$K=\bigcup_{k=1}^{m}L_{k}$
+and so $d$ is $C^{1}$ in $\Omega\setminus K$.
+We now use a standard cut-off argument.
+We denote by $d_{E}(\cdot,K)$ the Euclidean distance to the set $K$ and we consider smooth functions $\psi_{\epsilon}$, $\epsilon>0$, such that $0\leq\psi_{\epsilon}\leq 1$ and
+$\left\{\begin{array}[]{ll}\psi_{\epsilon}(\xi)=0,&\mbox{ if }d_{E}(\xi,K)<\epsilon,\\[2.84544pt] \psi_{\epsilon}(\xi)=1,&\mbox{ if }d_{E}(\xi,K)>2\epsilon,\\[2.84544pt] |\nabla\psi_{\epsilon}(\xi)|\leq\frac{c}{\epsilon},&\mbox{ for all }\xi\in{\mathbb{H}}.\end{array}\right.$
+We then define $\phi_{\epsilon}=\phi\,\psi_{\epsilon}$.
+Let $k\in\{1,\ldots,m\}$. Integrating by parts and using Proposition 6 we obtain
+$\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi=$ $-\int_{A_{k}}\phi_{\epsilon}\,\Delta_{H}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS$
+$\geq$ $\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
+where $\nu_{k,H}$ is defined as above relative to $A_{k}$. Adding over $k$ we arrive at
+(31) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi\geq\sum_{k=1}^{m}\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
 Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$.
-Let us fix such a set $A_{j}$, $j\neq k$.
-Denoting by $\nu_{H}$ the horizontal normal vector which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S$ from $A_{k}$ and $A_{j}$ add up to
+Let us fix $k,j$ such a set $A_{k}$ and $A_{j}$ share such a common boundary $S_{kj}$.
+Letting, as above, $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ where $\nu$ is the normal vector on $S_{kj}$ which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S_{kj}$ from $A_{k}$ and $A_{j}$ add up to
 […]
-The surface $S$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$. We therefore have
-$(\nabla_{{\mathbb{H}}^{n}}\,d_{k}-\nabla_{{\mathbb{H}}^{n}}\,d_{j})\cdot\nu_{H}$
+The surface $S_{kj}$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S_{kj}$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$.
+We therefore have on $S_{kj}$
+$(\nabla_{H}\,d_{k}-\nabla_{H}\,d_{j})\cdot\nu_{H}$
 $=\big{(}\nabla_{\!x}\,d_{k}-\nabla_{\!x}\,d_{j}+2d_{k,t}y-2d_{j,t}y,\nabla_{\!y}\,d_{k}-\nabla_{\!y}\,d_{j}-2d_{k,t}x+2d_{j,t}x\big{)}$
@@
 $=\big{(}\lambda\nu_{x}+2\lambda\nu_{t}y,\lambda\nu_{y}-2\lambda\nu_{t}x\big{)}\cdot\big{(}\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x\big{)}$
-$=\lambda\Big{(}|\nu_{x}+2\nu_{t}y|^{2}+|\nu_{y}-2\nu_{t}x|^{2}\Big{)}$
+$=\lambda\Big{(}(\nu_{x}+2\nu_{t}y)^{2}+(\nu_{y}-2\nu_{t}x)^{2}\Big{)}$
 $\geq 0.$
-Combining the above completes the proof of (i). Part (ii) is an immediate consequence of part (i) and Theorem 1. ∎
-As an immediate consequence of Theorem 10 we obtain the geometric uncertainty principle on the convex set $\Omega\subset{\mathbb{H}}^{n}$ with respect to the gauge pseudodistance on ${\mathbb{H}}^{n}$.
+Hence the LHS of (31) is non-negative.
+Noting that $\nabla\phi_{\epsilon}\to\nabla\phi$ in $L^{1}(\Omega)$ as $\epsilon\to 0$ completes the proof.
+Combining the above completes the proof of (i).
+Part (ii) is an immediate consequence of part (i) and Theorem 1.The optimality of the constant follows from Theorem 8.∎
+As an immediate consequence of Theorem 10 we obtain the geometric uncertainty principle on the convex set $\Omega\subset{\mathbb{H}}^{n}$ with respect to the gauge pseudo-distance on ${\mathbb{H}}^{n}$.
 Corollary 11.
-Let $D\subset\mathbb{H}^{n}$ be either a bounded convex polytope or an arbitary half-space in ${\mathbb{H}}^{n}$ and let
+Let $D\subset\mathbb{H}^{n}$ be either a bounded convex polytope or an arbitrary half-space in ${\mathbb{H}}^{n}$ and let
 $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial D)$
@@
 $\geq$ $\frac{1}{4}\left(\int_{D}u^{2}{\rm d}\xi\right)^{2}.$
+3.3. Hardy inequalities with respect to the Carnot-Carathéodory distance
+In this section we consider geometric Hardy inequalities in the Heisenberg group with respect to the Carnot-Carathéodory distance.
+In the case of the Heisenberg group ${\mathbb{H}}^{n}$, it has been proved in [BGG00, CCG07] (see also [DLZ24]) that the Carnot-Carathéodory distance of a point $\xi=(x,y,t)$ to the origin is given by
+(32) $\rho(\xi)=\left\{\begin{array}[]{ll}\frac{\phi}{\sin\phi}r,&\mbox{ if }(x,y)\neq(0,0),\\[5.69046pt] \sqrt{\pi|t|},&\mbox{ if }(x,y)=(0,0),\end{array}\right.$
+where $r=\sqrt{|x|^{2}+|y|^{2}}$ and the angle $\phi\in(0,\pi)$ is uniquely determined by the requirement
+(33) $\mu(\phi):=\frac{2\phi-\sin(2\phi)}{2\sin^{2}\phi}=\frac{t}{r^{2}}.$
+For a given domain $\Omega\subset\mathbb{H}^{n}$ the induced distance to the boundary is given by
+(34) $d_{\rho}(\xi)={\rm dist}_{\rho}(\xi,\partial\Omega)=\inf\{\rho((\xi^{\prime})^{-1}\xi),\;\xi^{\prime}\in\partial\Omega\},\qquad\xi\in\Omega.$
+As in Section 3.2, we first consider the case of the half-space
+$\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}.$
+By the above, it is then easy to see that for $\xi=(x,y,t)\in\Pi_{0}$ and $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ we have
+(35) $d_{\rho}(\xi,\xi^{\prime})=\left\{\begin{array}[]{ll}\frac{\phi}{\sin\phi}\sqrt{|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}},&(x^{\prime},y^{\prime})\neq(x,y),\\[5.69046pt] \sqrt{\pi t}\,,&(x^{\prime},y^{\prime})=(x,y),\end{array}\right.$
+where $\phi\in(0,\pi)$ is implicitly given by
+(36) $\mu(\phi)=\frac{t+2(x\cdot y^{\prime}-x^{\prime}\cdot y)}{|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}}.$
+We set
+$g(\phi)=\frac{\phi}{\sin\phi}$
+and define the function
+$F(x^{\prime},y^{\prime})=g(\phi)\sqrt{|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}},\qquad(x^{\prime},y^{\prime})\neq(x,y).$
+where $\phi\in(0,\pi)$ is determined by (36) with $\xi\in\Pi_{0}$ being fixed.
+Lemma 12.
+Let $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(0,0)$.
+The function $F(x^{\prime},y^{\prime})$ has a unique critical point $(x^{\prime},y^{\prime})$ in the set $\mathbb{R}^{2n}\setminus\{(x,y)\}$.
+Moreover this critical point is given by
+(37) $\left\{\begin{array}[]{l}(x_{i}^{\prime}-x_{i})\cot\phi=y_{i}\;,\\ (y_{i}^{\prime}-y_{i})\cot\phi=-x_{i}\;,\qquad i=1,\ldots,n,\end{array}\right.$
+where $\phi$ is the unique solution in $(0,\pi/2)$ of the equation
+(38) $\frac{2\phi+\sin(2\phi)}{2\cos^{2}\phi}=\frac{t}{r^{2}}.$
+Proof.
+For notational simplicity we set $X=(x,y)$ and $X^{\prime}=(x^{\prime},y^{\prime})\neq X$. We then have
+(39) $F_{x_{i}^{\prime}}(x^{\prime},y^{\prime})=g^{\prime}(\phi)\phi_{x_{i}^{\prime}}|X^{\prime}-X|+g(\phi)\frac{x_{i}^{\prime}-x_{i}}{|X^{\prime}-X|}.$
+Differentiating (36) we get
+$\mu^{\prime}(\phi)\phi_{x_{i}^{\prime}}=$ $-\frac{2y_{i}}{|X^{\prime}-X|^{2}}-\frac{2\big{(}t+2(x\cdot y^{\prime}-x^{\prime}\cdot y)\big{)})(x_{i}^{\prime}-x_{i})}{|X^{\prime}-X|^{4}}$
+$=$ $\frac{-2y_{i}-2\mu(\phi)\,(x_{i}^{\prime}-x_{i})}{|X^{\prime}-X|^{2}}.$
+Hence going back to (39) we see that $F_{x_{i}^{\prime}}(x^{\prime},y^{\prime})=0$ if and only if
+(40) $\Big{(}g(\phi)\mu^{\prime}(\phi)-2g^{\prime}(\phi)\mu(\phi)\Big{)}(x_{i}^{\prime}-x_{i})=2g^{\prime}(\phi)y_{i}.$
+But
+$\mu^{\prime}(\phi)=\frac{2(\sin\phi-\phi\cos\phi)}{\sin^{3}\phi},\qquad g^{\prime}(\phi)=\frac{\sin\phi-\phi\cos\phi}{\sin^{2}\phi}$
+and
+$g(\phi)\mu^{\prime}(\phi)-2g^{\prime}(\phi)\mu(\phi)=2\;\frac{\cos\phi(\sin\phi-t\cos\phi)}{\sin^{3}\phi}.$
+so (40) takes the form $(x_{i}^{\prime}-x_{i})\cot\phi=y_{i}$.
+Similar considerations show that $F_{y_{i}^{\prime}}(x^{\prime},y^{\prime})=0$ if and only if $(y_{i}^{\prime}-y_{i})\cot\phi=-x_{i}$.
+Hence we conclude that a point $(x^{\prime},y^{\prime})\neq(x,y)$ is a critical point of $F$ if and only if (37) is satisfied where $\phi\in(0,\pi)$ is given by (36).
+Assume now that $(x^{\prime},y^{\prime})\neq(x,y)$ is a critical point of the function $F$. From (37) we then have that $\phi\neq\pi/2$. Moreover, again from (37),
+$x_{i}(x_{i}^{\prime}-x_{i})+y_{i}(y_{i}^{\prime}-y_{i})=0\,,\qquad i=1,\ldots,n.$
+and
+$x\cdot y^{\prime}-x^{\prime}\cdot y=x\cdot(y^{\prime}-y)-(x^{\prime}-x)\cdot y=-r^{2}\tan\phi.$
+It then follows that
+(41) $|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}=r^{2}\tan^{2}\phi.$
+These together with (36) give
+$\mu(\phi)=\frac{t-2r^{2}\tan\phi}{(r^{2}\tan^{2}\phi}.$
+Recalling the definition of the function $\mu$, cf. (33), we conclude that $\phi$ satisfies (38). This procedure defines a map
+$\Pi_{0}\setminus\{(0,0,t):t>0\}\ni(x,y,t)\mapsto\phi=\phi(x,y,t)\in(0,\pi).$
+We shall write $\phi=\phi(r,t)$, $r=\sqrt{|x|^{2}+|y|^{2}}$.
+This map is clearly continuous and, as we have seen, we have $\phi(r,t)\neq\pi/2$ for all $r,t>0$.
+But, by (38), $\phi(r,t)\to 0$ as $t\to 0+$ for any fixed $r>0$, hence by continuity we obtain that in fact $\phi(r,t)\in(0,\pi/2)$ for all $r,t>0$.
+Hence given $(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(0,0)$, if a critical point $(x^{\prime},y^{\prime})\neq(x,y)$ exists then it is unique and it is determined by (38) and (37).
+To complete the proof we observe that the argument works both ways: if $\phi\in(0,\pi/2)$ is defined by (38) then the point $(x^{\prime},y^{\prime})$ defined by (37) is different from $(x,y)$ and is indeed a critical point of $F$.
+∎
+Lemma 13.
+The Carnot-Carathéodory distance of a point $\xi=(x,y,t)\in\Pi_{0}$ to the boundary $\partial\Pi_{0}$ is given by
+$d_{\rho}(\xi)=\left\{\begin{array}[]{ll}\frac{\phi}{\cos\phi}r,&\mbox{ if }(x,y)\neq(0,0),\\[5.69046pt] \sqrt{\frac{\pi t}{2}},&\mbox{ if }(x,y)=(0,0),\end{array}\right.$
+where $r=\sqrt{|x|^{2}+|y|^{2}}$ and in the first case $\phi\in(0,\pi/2)$ is uniquely determined by the requirement
+(42) $\frac{2\phi+\sin(2\phi)}{2\cos^{2}\phi}=\frac{t}{r^{2}}.$
+Moreover in case $(x,y)\neq(0,0)$ the distance is realized at a unique point $\xi^{\prime}\in\partial\Pi_{0}$ while in case $(x,y)=(0,0)$ it is realized at all points of the circle with center at the origin and radius $\sqrt{2t/\pi}$.
+Proof.
+(i) Case $(x,y)\neq(0,0)$.
+Let $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(0,0)$ be given and let $(x^{\prime},y^{\prime})\neq(x,y)$ be the critical point of $F$ determined in Lemma 12.
+Writing $\xi^{\prime}=(x^{\prime},y^{\prime},0)$ and recalling (35) and (41) we have
+$d_{\rho}(\xi,\xi^{\prime})=$ $F(x^{\prime},y^{\prime})$
+$=$ $\frac{\phi}{\sin\phi}\sqrt{|x^{\prime}-x|^{2}+|y^{\prime}-y|^{2}}$
+$=$ $\frac{\phi}{\sin\phi}r\;\tan\phi$
+(43) $=$ $\frac{\phi}{\cos\phi}r.$
+The proof will be complete if we prove that the distance of $\xi$ to the point $(x,y,0)$ is strictly larger than $d_{\rho}(\xi,\xi^{\prime})$. To see this we recall that, by (35),
+$d_{\rho}\big{(}\xi\,,\,(x,y,0)\big{)}=\sqrt{\pi t}.$
+Applying the elementary inequality
+$\phi^{2}<\frac{\pi}{2}\big{(}2\phi+\sin(2\phi)\big{)}\;,\qquad 0<\phi<\frac{\pi}{2}.$
+we thus have
+$d_{\rho}^{2}(\xi,\xi^{\prime})$ $=\frac{\phi^{2}}{\cos^{2}\phi}r^{2}$
+$<\frac{2\phi+\sin(2\phi)}{2\cos^{2}\phi}\pi r^{2}$
+$=\frac{t}{r^{2}}\,\pi r^{2}$
+$=d_{\rho}^{2}\big{(}\xi\,,\,(x,y,0)\big{)}.$
+(ii) Case $(x,y)=(0,0)$. Let $t>0$ be fixed. We recall that
+$d_{\rho}\big{(}(0,0,t)\,,(0,0,0)\big{)}=\sqrt{\pi t}.$
+Now, for $(x^{\prime},y^{\prime})\neq(0,0)$ we define
+$G(x^{\prime},y^{\prime})=d_{\rho}^{2}\big{(}(x^{\prime},y^{\prime},0)\,,\,(0,0,t)\big{)}=\frac{\phi^{2}}{\sin^{2}\phi}r^{\prime 2}$
+where $r^{\prime 2}=x^{\prime 2}+y^{\prime 2}$ and $\phi\in(0,\pi)$ is defined by
+$\frac{2\phi-\sin(2\phi)}{2\sin^{2}\phi}=\frac{t}{r^{\prime 2}}.$
+Hence
+$G(x^{\prime},y^{\prime})=\frac{2\phi^{2}}{2\phi-\sin(2\phi)}t.$
+The function $\phi\mapsto\phi^{2}/(2\phi-\sin(2\phi))$ is minimized for $\phi=\pi/2$ in which case it is equal to $\pi/2$.
+Hence $d_{\rho}(0,0,t)=\sqrt{\pi t/2}$.
+Moreover we have $\phi=\pi/2$ precisely for the points $(x^{\prime},y^{\prime})$ for which $r^{\prime 2}=2t/\pi$.
+This completes the proof.
+∎
+Proposition 14.
+Let $\Pi_{0}=\{(x,y,t)\in{\mathbb{H}}^{n},t>0\}$ and let $d_{\rho}=d_{\rho}(\xi)$, $\xi\in\Pi_{0}$, denote the corresponding Carnot-Carathéodory distance to the boundary $\partial\Pi_{0}$.
+Then for any fixed $r\neq 0$ we have
+$d_{\rho}(\xi)=\frac{t}{2r}+o(t)\,,\quad\mbox{as $t\rightarrow 0^{+}$. }$
+Proof.
+The proof follows using standard arguments by Lemma 13, since by (42) $t\rightarrow 0$ implies that $\phi\rightarrow 0$. ∎
+Theorem 15.
+Let $\Pi$ be an arbitrary half-space in ${\mathbb{H}}^{n}$ and let $d_{\rho}(\xi)$, $\xi\in{\mathbb{H}}^{n}$, denote the Carnot-Carathéodory distance to the boundary $\partial\Pi$. Then
+${\rm(i)}$ $\quad\Delta_{{\mathbb{H}}^{n}}d_{\rho}\leq 0\;\mbox{ in the distributional sense in }\Pi.$
+${\rm(ii)}$ For any $p>1$ the Hardy inequality
+$\hskip 42.67912pt\int_{\Pi}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Pi}\frac{|u|^{p}}{d_{\rho}^{p}}{\rm d}\xi\;,\quad u\in{C^{\infty}_{c}}(\Pi),$
+is valid. Moreover the constant is the best possible.
+Proof.
+For simplicity we write $d$ instead of $d_{\rho}$.
+$({\rm i})$ We first note that by invariance under group action we may assume that $\Pi=\Pi_{0}$.
+By Lemma 13 the distance to the boundary of a point $\xi=(x,y,t)$, $(x,y)\neq(0,0)$, is given by
+$d(r,t)=\frac{\phi}{\cos\phi}r=:B(\phi)r.$
+We will see below that $d_{r}(r,t)\leq 0$. Therefore, we have
+(44) $\Delta_{{\mathbb{H}}^{n}}d=$ $\;d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt}$
+$\leq$ $\;d_{rr}+\frac{1}{r}d_{r}+4r^{2}d_{tt}$
+$=$ $\;B^{\prime\prime}(\phi)\Big{(}r\phi_{r}^{2}+4r^{3}\phi_{t}^{2}\Big{)}+B^{\prime}(\phi)\Big{(}r\phi_{rr}+3\phi_{r}+4r^{3}\phi_{tt}\Big{)}+\frac{1}{r}B(\phi).$
+The various partial derivatives of $\phi=\phi(r,t)$ are computed from the relation
+(45) $Q(\phi):=\frac{2\phi+\sin(2\phi)}{2\cos^{2}\phi}=\frac{t}{r^{2}}.$
+Differentiating we find
+$\phi_{t}=\frac{1}{r^{2}Q^{\prime}(\phi)}\;,\qquad\phi_{tt}=-\frac{Q^{\prime\prime}(\phi)}{r^{4}\,(Q^{\prime}(\phi))^{3}}\;,\qquad\phi_{r}=-\frac{2Q(\phi)}{r\,Q^{\prime}(\phi)}\;,$
+$\phi_{rr}=\frac{Q(\phi)}{r^{2}(Q^{\prime}(\phi))^{3}}\Big{(}6\,(Q^{\prime}(\phi))^{2}-4Q^{\prime\prime}(\phi)\,Q(\phi)\Big{)}.$
+Substituting in (44) we arrive at
+(46) $\Delta_{{\mathbb{H}}^{n}}d\leq\frac{1}{r(Q^{\prime})^{3}}\bigg{(}4(Q^{2}+1)(B^{\prime\prime}Q^{\prime}-Q^{\prime\prime}B^{\prime})+B\,(Q^{\prime})^{3}\bigg{)}.$
+Let $A$ denote the term in large brackets in (46). We have
+$B^{\prime}(\phi)=\frac{\cos\phi+\phi\,\sin\phi}{\cos^{2}\phi}\;,\qquad\quad B^{\prime\prime}(\phi)=\frac{\phi+\phi\sin^{2}\phi+\sin(2\phi)}{\cos^{3}\phi},$
+$Q^{\prime}(\phi)=2\,\frac{\cos\phi+\phi\,\sin\phi}{\cos^{3}\phi}\;,\qquad\quad Q^{\prime\prime}(\phi)=2\,\frac{\phi+2\phi\sin^{2}\phi+3\sin(2\phi)}{\cos^{4}\phi}.$
+Therefore
+$d_{r}(r,t)=B(\phi)+B^{\prime}(\phi)\phi_{r}\,r=-\sin\phi\leq 0$
+as claimed. Substituting we find after some computations that
+$A=$ $-\frac{8}{\cos^{9}\phi}\cdot\bigg{(}(1+2\cos^{4}\phi-3\cos^{2}\phi)\phi^{3}+\cos\phi\sin\phi\,(3-5\cos^{2}\phi)\phi^{2}$
+$\qquad\qquad\qquad+\cos^{2}\phi\,(3-4\cos^{2}\phi)\phi+\cos^{3}\phi\sin\phi\bigg{)}.$
+Now, the term in the large brackets above can be written as
+$\phi^{3}\sin^{4}\phi+\phi^{2}\sin^{2}\phi\,\cos\phi\,(3\sin\phi-\phi\cos\phi)$
+$\qquad\quad+\phi\sin\phi\cos^{2}\phi\,(3\sin\phi-2\phi\cos\phi)+\cos^{3}\phi\,(\sin\phi-\phi\cos\phi)$
+Each of these four terms is non-negative for $\phi\in(0,\pi/2)$; hence $A\leq 0$ and (i) has been proved.
+(ii) The required Hardy inequality will follow by applying Theorem 1 provided we establish that $\Delta_{p,{\mathbb{H}}^{n}}d\leq 0$ in the distributional sense in $\Pi$.
+Now, by [MSC01, Theorem 3.1] (or by a direct computation for $r\neq 0$) we have $|\nabla_{{\mathbb{H}}^{n}}d|=1$ a.e.
+Hence the above condition can be simplified to $\Delta_{{\mathbb{H}}^{n}}d\leq 0$ which is in particular independent of $p>1$; see also [BFT04].
+So the result follows from (i).
+The sharpness of the constant follows from Theorem 1 (b), Proposition 14 and the fact that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e., cf.
+[MSC01].
+∎
+The next proposition provides a more detailed picture concerning nearest boundary points. We denote $r^{\prime 2}=|x^{\prime}|^{2}+|y^{\prime}|^{2}$.
+Proposition 16.
+$(1a)$ Any point $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(0,0)$ has a unique nearest boundary point $(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$.
+Moreover $(x^{\prime},y^{\prime})$ is different from $(x,y)$ and from $(0,0)$ and there holds $d_{\rho}(\xi)<\pi r^{\prime}/2$.
+$(1b)$ Let $t>0$. The point $(0,0,t)$ has as nearest boundary points all points $(x^{\prime},y^{\prime},0)$ with $x^{\prime 2}+y^{\prime 2}=2t/\pi$.
+$(2a)$ Conversely, let $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ with $(x^{\prime},y^{\prime})\neq(0,0)$ and $\rho>0$ be given.
+If $\rho<\pi r^{\prime}/2$ then there are exactly two points $\xi=(x,y,t)\in\Pi_{0}$ whose nearest boundary point is $\xi^{\prime}$ and for which $d_{\rho}(\xi)=\rho$.
+Moreover exactly one of these points lies on the $t$-axis.
+If $\rho\geq\pi r^{\prime}/2$ then there is only one such point and it lies on the $t$-axis.
+$(2b)$ No point in $\Pi_{0}$ has $(0,0,0)$ as its nearest boundary point.
+Proof.
+(1a) We have already seen in the proof of Theorem 13 that given $\xi\in\Pi_{0}$ the nearest boundary point $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ is uniquely defined.
+Moreover, setting $r^{\prime 2}=x^{\prime 2}+y^{\prime 2}$ we have from (37) that
+$r^{\prime 2}=(1+\tan^{2}\phi)\,r^{2}=\frac{r^{2}}{\cos^{2}\phi}$
+where $\phi\in(0,\pi/2)$ is given by (42). Hence
+(47) $d_{\rho}(\xi)=\frac{\phi}{\cos\phi}r=\phi r^{\prime}<\frac{\pi r^{\prime}}{2}.$
+(1b) This is contained in Theorem 13.
+(2a) Assume that $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(x^{\prime},y^{\prime})$ has $\xi^{\prime}$ as its nearest boundary point and that $d(\xi)=\rho$.
+By (47) $\phi=\rho/r^{\prime}$ and the point $\xi$ is now uniquely determined by (37) and (42).
+It is now easy to see that this point $\xi$ has indeed $\xi^{\prime}$ as its nearest boundary point and $d_{\rho}(\xi)=\rho$.
+Moreover, by Theorem 13, the point $(0,0,2\rho^{2}/\pi)$ also has $(x^{\prime},y^{\prime},0)$ as one of its nearest boundary points and its distance to the boundary is $\rho$.
+In case $\rho\geq\pi r^{\prime}/2$ the first of these two points is not defined.
+Hence (2a) has been proved.
+(2b) We note that in case (1a) the nearest boundary point $(x^{\prime},y^{\prime},0)$ is not the origin since $(0,0)$ is not a critical point of the function $F$ in Lemma 12.
+In case (1b) the nearest boundary points are also different from the origin.
+Lemma 17.
+Let $\Omega\subset\mathbb{H}^{n}$ be a bounded convex polytope.
+The corresponding Carnot-Carathéodory distance to the boundary satisfies $\Delta_{{\mathbb{H}}^{n}}d_{\rho}\leq 0$ in the distributional sense in $\Omega$.
+Proof.
+The proof follows exactly the lines of the proof of Theorem 10 using Lemma 13 and Theorem 15 instead. ∎
+Theorem 18.
+Let $\Omega\subset\mathbb{H}^{n}$ be bounded and convex and let $d_{\rho}(\xi)$, $\xi\in\Omega$, denote the corresponding Carnot-Carathéodory distance to the boundary.
+Then for any $p>1$ the Hardy inequality
+$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega}\frac{|u|^{p}}{d_{\rho}^{p}}{\rm d}\xi\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
+is valid.
+Proof.
+For simplicity we write $d$ instead of $d_{\rho}$. Let $u\in{C^{\infty}_{c}}(\Omega)$ be fixed. We consider a bounded convex polytope $\Omega^{\prime}$ such that
+${\rm supp}(u)\subset\subset\Omega^{\prime}\subset\subset\Omega.$
+Let us denote by $d^{\prime}$ the Carnot-Carathéodory distance to $\partial\Omega^{\prime}$. Combining Theorem 1 and Lemma 17 we obtain that
+$\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d^{\prime p}}{\rm d}\xi\;.$
+Hence, since $d^{\prime}\leq d$ in $\Omega^{\prime}$,
+$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi=\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d^{\prime p}}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega}\frac{|u|^{p}}{d^{p}}{\rm d}\xi$
+and the proof is complete. ∎
+Remark 4.
+We point that the fact that the key property that $d^{\prime}\leq d$, for $d$ and $d^{\prime}$ as in the proof of Theorem 18 in the bounded convex polytope $\Omega^{\prime}$ and reflects the geometric nature of the Carnot-Carathéodory distance, meaning in particular that that the latter is a “true” distance respecting the geometry of ${\mathbb{H}}^{n}$.
+The proof of the following corollary, which is the geometric uncertainty principle with respect to the Carnot-Carathéodory distance, follows the lines of Corollary 11 in the case of the gauge pseudodistance.
+Corollary 19.
+Let $D\subset\mathbb{H}^{n}$ be either a bounded convex domain or an arbitrary half-space in ${\mathbb{H}}^{n}$ and let
+$d_{\rho}(\xi)={\rm dist}_{\rho}(\xi,\partial D)$
+denote the corresponding Carnot-Carathéodory distance of $\xi\in D$ to the boundary $\partial D$. Then for any $u\in C_{c}^{\infty}(D)$ we have
+$\left(\int_{D}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}\xi\right)^{\frac{1}{2}}\left(\int_{D}d_{\rho}^{2}u^{2}{\rm d}\xi\right)^{\frac{1}{2}}\geq\frac{1}{2}\int_{D}u^{2}{\rm d}\xi\,.$
 4. Hardy inequalities on stratified groups of step two

### v3

@@
 In all cases the constant $((p-1)/p)^{p}$ is obtained.
-In the more general context of a stratified Lie group of step two we study the superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an alternative proof for the $L^{2}$-Hardy inequality on convex domains.
+In the more general context of a stratified Lie group of step two we study the $p$-superharmonicity and the weak $H$-concavity of the Euclidean distance to the boundary, thus obtaining an $L^{p}$-Hardy inequality on bounded convex domains which generalises previous results.
 Key words and phrases:
@@
 The literature on geometric Hardy inequalities in Euclidean space is large and we refer the interested reader to the works [Ba24, BEL15, Dav98, RS19] which provide an overview of the topic.
-On the other hand, subelliptic Hardy inequalities have been studied for quite a long time and the work [GL90] of Garofallo and Lanconelli in the 90’s opened up the research in this direction.
+On the other hand, subelliptic Hardy inequalities have been studied for quite a long time and the work [GL90] of Garofalo and Lanconelli in the 90’s opened up the research in this direction.
 By subelliptic Hardy inequalities, we mean Hardy-type inequalities considered in the setting of homogeneous Lie groups, and in particular stratified groups.
@@
 where $d(x)$, $x\in\Omega$, stands for the Euclidean distance to $\partial\Omega$ and the constant $1/4$ is the best possible.
-Later on in [RSS20] Ruzhansky et al.
-proved that if $\Omega$ is a half-space in any stratified group then for any $p>1$ there holds
-$\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}x\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{T}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
-where $\nabla_{H}$ denotes the horizontal gradient on $H$. For other results in this direction see also [Rus18].
+Later on in [RSS20b] Ruzhansky et al.
+proved that if $\Omega$ is a half-space in $\mathbb{H}^{n}$ and $p>1$ then
+$\int_{\Omega}|\nabla_{\mathbb{H}^{n}}u|^{p}{\rm d}x\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|\nabla_{\mathbb{H}^{n}}d|^{p}}{d^{p}}|u|^{p}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(\Omega).$
+A little earlier the same authors had showed that if $\Omega$ is a convex domain in any stratified group then
+$\int_{\Omega}|\nabla_{H}u|^{2}\,{\rm d}x\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{H}d|^{2}}{d^{2}}u^{2}\,{\rm d}x\,$,
+where $\nabla_{H}$ denotes the horizontal gradient [RSS20a].
+For other results in this direction see also [Rus18]; see also [PRS19] for an alternative approach based on a sub-Riemannian version of the Santaló formula.
 Our main interest in this work is to prove subelliptic geometric Hardy inequalities with best constant on the Heisenberg group ${\mathbb{H}}^{n}$, but also on any stratified group of step two.
@@
 The general result is then implemented in different contexts.
-In the case of the Euclidean distance on the Heisenberg group ${\mathbb{H}}^{n}$ we prove the following result which goes beyond the convexity condition.
+Let us first consider some known results about Hardy inequalities in Euclidean domains. Assume that $U\subset\mathbb{R}^{d}$ is a bounded domain with $C^{2}$ boundary and let $p>1$.
+One approach to study the Hardy inequality in $U$ is by using the fact that if we consider functions supported in a small neighbourhood of a boundary point, then the Hardy constant is $((p-1)/p)^{p}$.
+However, if we try to pass from this local inequality to a global one on $U$, then one ends up with the inequality
+$\int_{U}|\nabla u|^{p}{\rm d}x+c\int_{U}|u|^{p}{\rm d}x\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{U}\frac{|u|^{p}}{d^{p}}{\rm d}x\;,\qquad u\in{C^{\infty}_{c}}(U).$
+for some $c>0$ depending on $U$. It is not possible to avoid the $L^{p}$ term unless some geometric assumption is made on $U$.
+Indeed, for $p=2$, [Dav95] Davies constructed certain planar sectors with Hardy constant smaller that $1/4$.
+By approximating such a sector by an increasing sequence of bounded smooth domains one concludes the existence of bounded smooth domains with Hardy constant smaller than $1/4$.
+In the same article and for dimension $d\geq 3$, Davies constructed conical domains with Hardy constant arbitrarily small; again, this shows the existence of bounded smooth domains with the same property.
+Concerning the general case $p>1$, elemantary considarations show that when the domain is the complement of a ball in $\mathbb{R}^{d}$ and $p>(d+1)/2$ then the (global) Hardy constant is smaller than $((p-1)/p)^{p}$.
+A simple argument then shows the existence of tori for which the same holds true.
+It is a central question to find conditions on the domain $U$ under which the Hardy constant is equal to $((p-1)/p)^{p}$.
+In the context of the Heisenberg group $\mathbb{H}^{n}$ one similarly expects that in order to obtain the Hardy constant $((p-1)/p)^{p}$ some geometric condition is necessary on the domain.
+Indeed, Proposition 3 below uses a lifting argument to show that each of the above examples for Euclidean domains induces a corresponding example in $\mathbb{H}^{n}$ for the Hardy inequality with Euclidean distance.
+The main idea here is to lift the Euclidean domain $U\subset\mathbb{R}^{n}$ to the domain $U\times\mathbb{R}\subset{\mathbb{H}}^{n}$ in the Heisenberg group.
+The condition introduced in this article, namely the $p$-horizontal superharmonicity of the distance function, turns out to be a natural condition.
+For the case of the Euclidean distance in a Carnot group of step two it applies to any convex domain for any $p>1$, see Section 4.
+But it is also valid to domains that are not necessarily convex as is seen in Section 2.
+Moreover it is applied, as already mentioned, to the Carnot-Carathéodory distance and the gauge pseudodistance.
+We now present the main results of this article.
+The torus $T\subset\mathbb{H}^{n}$ defined in the next theorem is not a convex domain but satisfies the geometric assumption $\Delta_{p,\mathbb{H}^{n}}d\leq 0$ for any $p>1$.
 The precise value of the constant $\beta(p,n)$ is given in Proposition 4.
@@
 then
+$\Delta_{p,\mathbb{H}^{n}}\,d\leq 0\,\quad\text{in the weak sense in}\,\,T\,$,
+and
 $\int_{T}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{T}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{p}}{d^{p}}|u|^{p}{\rm d}\xi\;,\qquad u\in{C^{\infty}_{c}}(T).$
+Moreover the constant $((p-1)/p)^{p}$ is sharp.
 Another class of distances in a stratified group consists of those induced by a homogenous quasi-norm.
@@
 Moreover the constant is the best possible in both cases.
-Denoting by $d_{\rho}$ the Carnot-Carathéodory distance to the boundary, we obtain the following
+Denoting by $d_{\rho}$ the Carnot-Carathéodory distance (control distance) to the boundary, we obtain the following
 Theorem C. Let $p>1$ and let $D\subset{\mathbb{H}}^{n}$ be a bounded and convex domain or a half-space in ${\mathbb{H}}^{n}$. There holds
@@
 In the case of the half-space the constant is sharp.
-The appereance of the factor $|\nabla_{{\mathbb{H}}^{n}}d_{N}|$ in Theorem B – while a similar factor does not appear in Theorem C – should be seen in the light of a result of [MSC01] a special case of which states that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e.
+The appearance of the factor $|\nabla_{{\mathbb{H}}^{n}}d_{N}|$ in Theorem B – while a similar factor does not appear in Theorem C – should be seen in the light of a result of [MSC01] a special case of which states that $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e.
 A more interesting difference between the above two homogeneous norms relates to nearest boundary points in the case of a subspace $\Pi\subset{\mathbb{H}}^{n}$.
@@
 where $s={\rm dist}(\xi,\partial\Pi)=d(\xi,\xi^{\prime})$ is different in the two cases.
-In the case of the quasi-norm (1) the above is a simple 1-1 correspondance, as in the Euclidean case, cf.
+In the case of the quasi-norm (1) the above is a simple 1-1 correspondence, as in the Euclidean case, cf.
 Proposition 10.
@@
 Rickly prove that in the case of the Heisenberg group $\mathbb{H}$, if a set is geodesically convex and contains at least three points that do not lie on the same geodesic, then it necessarily coincides with $\mathbb{H}$.
-In our context the relevant notions of convexity of sets and functions are the ones introduced at the same time by Lu, Manfredi and Stroffolini in [LMS03] on the Heisenberg group and by Danielli, Garofallo and Nhieu in [DGN03] on any stratified group.
+In our context the relevant notions of convexity of sets and functions are the ones introduced at the same time by Lu, Manfredi and Stroffolini in [LMS03] on the Heisenberg group and by Danielli, Garofalo and Nhieu in [DGN03] on any stratified group.
 These notions are the analogues of the corresponding ones in the abelian case $\mathbb{R}^{n}$ but with a twist; the condition refers to a convex combination of two elements $g,g^{\prime}$ for which, additionally, $g^{\prime}\in H_{g}$, i.e.
 $g^{\prime}$ lies in the horizontal plane passing through $g$; see Section 4.1 for the precise definitions.
-Exploring properties of the so-called weakly $H$-concave functions (see Definition 20) and using a result from [DGN03] we prove the following theorem, part (iii) of which is contained in [RSS20].
+Exploring properties of the so-called weakly $H$-concave functions (see Definition 21) and using a result from [DGN03] we prove the following theorem, part (iii) of which generalizes [RSS20b, Corollary 2.2].
 Theorem D. Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
 […]
-${\rm(ii)}$ $\quad\Delta_{{\mathbb{H}}^{n}}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
+${\rm(ii)}$ $\quad\Delta_{p,H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
 ${\rm(iii)}$ The Hardy inequality
-$\hskip 42.67912pt\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{2}{\rm d}g\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{{\mathbb{H}}^{n}}d|^{2}}{d^{2}}u^{2}{\rm d}g\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
+$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}g\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}g\;,\qquad u\in{C^{\infty}_{c}}(\Omega),$
 is valid.
@@
 where
-$[V_{i},V_{j}]={\rm span}\{[X,Y]:\;X\in V_{i},\;Y\in V_{j}\}.$
+$V_{i},V_{j}]={\rm span}\{[X,Y]:\;X\in V_{i},\;Y\in V_{j}\}.$
 Such a stratification naturally equips $G$ with a non-anisotropic dilation structure $\delta_{\lambda}:G\rightarrow G$, $\lambda>0$, and makes $G$ a homogeneous Lie group.
@@
 In what follows, we will say that a function is CC-Lipschitz if it is Lipschitz with respect to the Carnot-Carathéodory distance (equivalently, with respect to the distance induced by any homogeneous quasi-norm).
-Part (a) of the next theorem is essentially contained in [RSS20] but we include the short proof of it because of its central role in the present article.
+Part (a) of the next theorem is essentially contained in [RSS20b] but we include the short proof of it because of its central role in the present article.
 Theorem 1.
@@
 We then have
-$d(g)-d(g^{\prime})$$\leq$ lim $\liminf\big{[}N(g^{-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
-$=$ lim $\liminf\big{[}N(g^{-1}g^{\prime}g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
-$\leq$ lim $\liminf\big{[}\beta N(g^{-1}g^{\prime})+N(g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big{]}$
+$d(g)-d(g^{\prime})$ $\leq$ $\liminf\big[N(g^{-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big]$
+$=$ $\liminf\big[N(g^{-1}g^{\prime}g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big]$
+$\leq$ $\liminf\big[\beta N(g^{-1}g^{\prime})+N(g^{\prime-1}b_{n}^{\prime})-N(g^{\prime-1}b_{n}^{\prime})\big]$
 $=$ $\beta N(g^{-1}g^{\prime})\,.$
 Similarly we can show that
-$d(g^{\prime})-d(g)\leq C\beta\,d_{N}g^{\prime},g)\,,$
+$d(g^{\prime})-d(g)\leq C\beta\,d_{N}(g^{\prime},g)\,$,
 and (4) follows. ∎
@@
 In the section we consider geometric Hardy inequalities on the Heisenberg group $\mathbb{H}^{n}$.
-In the first part we consider the Euclidean distance and prove the validity of the Hardy inequality with best constant on certain torii under suitable assumptions on the radii.
+In the first part we consider the Euclidean distance and prove the validity of the Hardy inequality with best constant on certain tori under suitable assumptions on the radii.
 In the second part we study the geometric Hardy inequality for the pseudodistance induced by the quasi-norm (1).
@@
 $X_{i}=\partial_{x_{i}}+2y_{i}\partial_{t}\;,\qquad Y_{i}=\partial_{y_{i}}-2x_{i}\partial_{t}\,,\quad i=1,\ldots,n,$
-form the canonical basis basis of the first stratum and the associated horizontal gradient and horizontal Laplacian on ${\mathbb{H}}^{n}$ are given respectively by
+form the canonical basis of the first stratum and the associated horizontal gradient and horizontal Laplacian on ${\mathbb{H}}^{n}$ are given respectively by
 $\nabla_{{\mathbb{H}}^{n}}=(X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}),$
@@
 So in the current section we denote the horizontal gradient and Laplacian by $\nabla_{{\mathbb{H}}^{n}}$ and $\Delta_{{\mathbb{H}}^{n}}$, respectively, to emphasize that the obtained results refer to the particular case of ${\mathbb{H}}^{n}$.
-3.1. Hardy inequalities with respect to the Euclidean distance on a torus
-We will see here that Theorem 1 can be applied in the case of the Euclidean distance on a torus and thus goes beyond the convexity assumption of [Lar16].
-We first note that for any $u\in C^{2}({\mathbb{H}}^{n})$ we have
+3.1. Hardy inequalities with respect to the Euclidean distance: remarks and counterexamples
+In this section we consider geometric Hardy inequalities in domains $\Omega\subset\mathbb{H}^{n}$ with respect to the Euclidean distance.
+The next proposition, when combined with the examples of Euclidean domains for which the Hardy constant is smaller than $((p-1)/p)^{p}$, shows the one cannot expect that a bounded $C^{2}$ domain in $\mathbb{H}^{n}$ has constant $((p-1)/p)^{p}$ unless some geometric condition is imposed.
+Proposition 3.
+Let $U\subset\mathbb{R}^{2n}$ be a domain. The Hardy constant of $\Omega=U\times\mathbb{R}$ as a domain in $\mathbb{H}^{n}$ cannot exceed the Hardy constant of $U$.
+Proof.
+Let $v\in C^{\infty}_{c}(U)$ and $\psi\in{C^{\infty}_{c}}(\mathbb{R})$ be given. We define
+$u(x,y,t)=v(x,y)\psi(t)\;,\qquad(x,y)\in U,\;t\in\mathbb{R},$
+A simple computation gives that for any $\epsilon>0$,
+$|\nabla_{\mathbb{H}^{n}}u|^{2}$ $=|\nabla v|^{2}\psi^{2}+4r^{2}v^{2}(\psi^{\prime})^{2}+4v\psi\psi^{\prime}\sum_{k=1}^{n}(y_{k}v_{x_{k}}-x_{k}v_{y_{k}})$
+$\leq(1+\epsilon)|\nabla v|^{2}\psi^{2}+4(1+\frac{1}{\epsilon})r^{2}v^{2}(\psi^{\prime})^{2}.$
+Combining this with the elementary inequality
+$|a+b|^{q}\leq(1+\epsilon)|a|^{q}+k_{\epsilon}|b|^{q}\;,\quad\epsilon>0,$
+we obtain that for any $\epsilon>0$ there exists $c_{\epsilon}>0$ such that
+$|\nabla_{\mathbb{H}^{n}}u|^{p}\leq(1+\epsilon)|\nabla v|^{p}|\psi|^{p}+c_{\epsilon}r^{p}|v|^{p}|\psi^{\prime}|^{p}\;,\quad\mbox{ in }U^{*}.$
+Since $d_{\Omega}(x,y,t)=d_{U}(x,y)$, we conclude that
+$\frac{\int_{\Omega}|\nabla_{\mathbb{H}^{n}}u|^{p}{\rm d}\xi}{\int_{\Omega}\frac{|u|^{p}}{d^{p}}{\rm d}\xi}\leq(1+\epsilon)\frac{\int_{U}|\nabla v|^{p}dx\,dy}{{\int_{U}\frac{|v|^{p}}{d^{p}}dx\,dy}}+c_{\epsilon}\frac{\int_{U}r^{p}|v|^{p}dx\,dy}{\int_{U}|v|^{p}dx\,dy}\cdot\frac{\int_{-\infty}^{\infty}|\psi^{\prime}|^{p}dt\,dy}{\int_{-\infty}^{\infty}|\psi|^{p}dt}.$
+The result follows by noting that the quotient involving $\psi$ can be made arbitrarily small. ∎
+We shall now consider the case of the torus $T$ which is not a convex domain and has been discussed in Theorem A. For $u\in C^{2}({\mathbb{H}}^{n})$ we have
 (5) $\Delta_{{\mathbb{H}}^{n}}u=\sum_{i=1}^{n}(u_{x_{i}x_{i}}+u_{y_{i}y_{i}})+4(|x|^{2}+|y|^{2})u_{tt}-2\sum_{i=1}^{n}(y_{i}u_{x_{i}t}-x_{i}u_{y_{i}t})$
@@
 $d(\xi)=\rho-\sqrt{(r-R)^{2}+t^{2}},\qquad\xi\in T,$
-and is smooth in $T$ except on the $(2n-1)$-dimensional ‘circle’
+and is smooth in $T$ except on the (2n1)(2n-1)-dimensional ‘circle’
 $S=\{\xi=(r,\omega,t):\;t=0,\;r=R\}.$
@@
 $\Delta_{\mathbb{H}^{n}}d=$ $-\big((r-R)^{2}+t^{2}\big)^{-\frac{3}{2}}\frac{1}{r}$
-(8) ×$\times\bigg{\{}\Big{[}(2n-1)(r-R)+4r^{3}\Big{]}(r-R)^{2}+\Big{[}r+(2n-1)(r-R)\Big{]}t^{2}\bigg{\}}.$
+(8) $\times\bigg\{\Big[(2n-1)(r-R)+4r^{3}\Big](r-R)^{2}+\Big[r+(2n-1)(r-R)\Big]t^{2}\bigg\}.$
 Proposition 4.
@@
 or equivalently
-(11) $016(p-1)r^{4}-\Big{(}8(p-2)+4a(2n+p-3)\Big{)}r^{2}+p-1-a(2n-1)\geq 0\,,\;\quad r>$ 0 .
+(11) $16(p-1)r^{4}-\Big(8(p-2)+4a(2n+p-3)\Big)r^{2}+p-1-a(2n-1)\geq 0\,,\;\quad r>0.$
 If such an $a$ has been found then we shall have $C_{1}\geq 0$ provided the radii of $T$ satisfy
@@
 At this point we need to distinguish different cases.
-Case $21<p<$ 2. In this case we choose
+Case $1<p<2$. In this case we choose
 $a=a_{1}(p,n):=\min\Big\{\frac{p-1}{2n-1},\;\frac{2(2-p)}{2n+p-3}\Big\}$
@@
 and $\beta(p,n)$ is given by
-[…]
+$\beta(p,n)=\max\Big\{2n+p-2\,,\;1+\frac{1}{a(p,n)}\Big\}.$
 Finally we to note that for $p>2$,
@@
 Remark 1.
-For $21<p<$ 2 we have that
+For $1<p<2$ we have that
 $\frac{2n-p+1}{2(2-p)}\geq 2n+p-2\quad\mbox{ iff }\quad\frac{3}{2}\leq p<2.$
-Moreover if we define $p_{0}=\sqrt{9n^{2}-8n+2}-3(n-1)$ then 3/$23/2<p_{0}<$ 2 and for $21<p<$ 2 we have
+Moreover if we define $p_{0}=\sqrt{9n^{2}-8n+2}-3(n-1)$ then $3/2<p_{0}<2$ and for $1<p<2$ we have
 $\max\Big\{\frac{2n+p-2}{p-1}\,,\;\frac{2n-p+1}{2(2-p)}\Big\}={\left\{\begin{array}[]{ll}{\frac{2n+p-2}{p-1},}&{\mbox{ if }1<p\leq p_{0},}\\ {\frac{2n-p+1}{2(2-p)},}&{\mbox{ if }p_{0}\leq p<2.}\end{array}\right.}$
@@
 Remark 2.
-It is evident from the argument in the above proof that if $\Omega\subset\mathbb{H}^{n}$ is a domain with $C^{2}$ boundary then the corresponding Hardy constant connot be larger than $((p-1)/p)^{p}$, provided there exists a point $\xi_{0}\in\partial\Omega$ such that $\nabla_{\mathbb{H}^{n}}d(\xi_{0})\neq 0$.
-This, of course, is very generic.
-We do not pursue this any further here, but we make two comments.
-1. Assume that $\Omega$ possesses cylindrical symmetry so that $d=d(r,t)$. Suppose that at a point $\xi_{0}\in\partial\Omega$ we have $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, that is
-$d_{x_{i}}+2y_{i}d_{t}=0\;,\qquad d_{y_{i}}-2x_{i}d_{t}=0\;,\qquad i=1,\ldots,n.$
-Multiplying by $x_{i}$ and $y_{i}$ respectively and adding we obtain
-$\sum_{i=1}^{n}(x_{i}d_{x_{i}}+y_{i}d_{y_{i}})=0,\qquad\mbox{ at the point }\xi_{0}.$
-Using cylindrical coordinates $(r,\omega,t)$ we then obtain that $d_{r}(\xi_{0})=0$. By cylindrical symmetry we have
-$d_{r}^{2}+d_{t}^{2}=|\nabla d|^{2}=1\;,\qquad\mbox{ in }T,$
-hence $d_{t}(\xi_{0})=1$ and therefore the tangent hyperplane at $\xi_{0}$ must be parallel to the hyperplane $\{t=0\}$. Hence generically we have $\nabla_{\mathbb{H}^{n}}d\neq 0$ on $\partial\Omega$
-2. Let $\Omega$ be a domain which does not necessarily possess some kind of symmetry. If $\xi_{0}\in\partial\Omega$ is a boundary point with $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, then we have
+It is evident from the argument in the above proof that if $\Omega\subset\mathbb{H}^{n}$ is a domain with $C^{2}$ boundary then the corresponding Hardy constant cannot be larger than $((p-1)/p)^{p}$, provided there exists a point $\xi_{0}\in\partial\Omega$ such that $\nabla_{\mathbb{H}^{n}}d(\xi_{0})\neq 0$.
+This is very generic.
+We refer to [M06] and references therein for results concerning the negligibility of the set of such points in a much wider context.
+In the case of the Heisenberg group $\mathbb{H}^{n}$ we can give a direct proof of this negligibility.
+Indeed, if $\xi_{0}\in\partial\Omega$ is a boundary point with $\nabla_{\mathbb{H}^{n}}d(\xi_{0})=0$, then we have
 $d_{x_{i}}^{2}=4y_{i}^{2}d_{t}^{2}\;,\qquad d_{y_{i}}^{2}=4x_{i}^{2}d_{t}^{2}\;,\quad i=1,\ldots,n,\quad\mbox{ at the point }\xi_{0}.$
@@
 denote the corresponding pseudodistance to the boundary. Then, for any $\xi=(x,y,t)\in\Pi_{0}$, $d_{N}(\xi)$ depends only on $r=\sqrt{|x|^{2}+|y|^{2}}$ and $t>0$. More precisely, we have
-$d_{N}(\xi)=d_{N}(r,t)={\left\{\begin{array}[]{ll}{\big{(}2r^{4}s^{2}-3tr^{2}s+t^{2}\big{)}^{\frac{1}{4}},}&{r>0,\;t>0,}\\ {t^{\frac{1}{2}},}&{r=0,\;t>0,}\end{array}\right.}$
+(16) $d_{N}(\xi)=d_{N}(r,t)={\left\{\begin{array}[]{ll}{\big(2r^{4}s^{2}-3tr^{2}s+t^{2}\big)^{\frac{1}{4}},}&{r>0,\;t>0,}\\ {t^{\frac{1}{2}},}&{r=0,\;t>0,}\end{array}\right.}$
 where for fixed $r,t>0$ the real number $s\in\mathbb{R}$ is the unique solution of the equation
@@
 Proof.
-For simplicity we will write $d$ instead of $d_{N}$. By group action (see also [Lar16, p340]) it is enough to consider the case $\Pi=\Pi_{0}$. Also, it is preferable to work with the function
-$G(r,t)=d(r,t)^{4}$
-instead of $d(r,t)$. To compute the various derivatives of $G(r,t)$ we recall from Lemma 16 that
+By group action (see also [Lar16, p340]) it is enough to consider the case $\Pi=\Pi_{0}$. Also, it is preferable to work with the function
+$G(r,t)=d_{N}(r,t)^{4}$
+instead of $d_{N}(r,t)$. To compute the various derivatives of $G(r,t)$ we recall from Lemma 17 that
 (21) $G(r,t)=2r^{4}s^{2}-3tr^{2}s+t^{2}=(t-r^{2}s)(t-2r^{2}s)\quad\mbox{ in }\Pi_{0}\setminus\{r=0\},$
@@
 $G_{tt}$ $=\frac{6s^{2}}{3s^{2}+2},\qquad\quad G_{rt}=-\frac{16rs^{3}}{3s^{2}+2}.$
-Setting $A=|\nabla_{\mathbb{H}^{n}}d|^{2}$ we have, cf. (9),
-(25) $\Delta_{p,\mathbb{H}^{n}}d=A^{\frac{p-4}{2}}\Big{(}A\,\Delta_{\mathbb{H}^{n}}d+\frac{p-2}{2}\big{(}d_{r}A_{r}+4r^{2}d_{t}A_{t}\big{)}\Big{)}.$
+Setting $A=|\nabla_{\mathbb{H}^{n}}d_{N}|^{2}$ we have, cf. (9),
+(26) $\Delta_{p,\mathbb{H}^{n}}d_{N}=A^{\frac{p-4}{2}}\Big(A\,\Delta_{\mathbb{H}^{n}}d_{N}+\frac{p-2}{2}\big(d_{N,r}A_{r}+4r^{2}d_{N,t}A_{t}\big)\Big).$
 We have
-$d_{r}=\frac{1}{4}G^{-\frac{3}{4}}G_{r},\qquad d_{t}=\frac{1}{4}G^{-\frac{3}{4}}G_{t},\qquad d_{rr}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{rr}$
+$d_{N,r}=\frac{1}{4}G^{-\frac{3}{4}}G_{r},\qquad d_{N,t}=\frac{1}{4}G^{-\frac{3}{4}}G_{t},\qquad d_{N,rr}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{rr}$
 and
-$d_{tt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{t}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{tt},\qquad d_{rt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}G_{t}+\frac{1}{4}G^{-\frac{3}{4}}G_{rt}.$
+$d_{N,tt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{t}^{2}+\frac{1}{4}G^{-\frac{3}{4}}G_{tt},\qquad d_{N,rt}=-\frac{3}{16}G^{-\frac{7}{4}}G_{r}G_{t}+\frac{1}{4}G^{-\frac{3}{4}}G_{rt}.$
 Moreover
-$A=d_{r}^{2}+4r^{2}d_{t}^{2}=\frac{1}{16}G^{-\frac{3}{2}}(G_{r}^{2}+4r^{2}G_{t}^{2})=G^{-\frac{3}{2}}r^{6}s^{6}(s^{2}+1),$
-$A_{r}$ $=2d_{r}d_{rr}+8rd_{t}^{2}+8r^{2}d_{t}d_{rt}$
+$A=d_{N,r}^{2}+4r^{2}d_{N,t}^{2}=\frac{1}{16}G^{-\frac{3}{2}}(G_{r}^{2}+4r^{2}G_{t}^{2})=G^{-\frac{3}{2}}r^{6}s^{6}(s^{2}+1),$
+$A_{r}$ $=2d_{N,r}d_{N,rr}+8rd_{t}^{2}+8r^{2}d_{N,t}d_{N,rt}$
 $=G^{-\frac{5}{2}}\Big(-\frac{3}{32}G_{r}^{3}+\frac{1}{8}G\,G_{r}G_{rr}+\frac{r}{2}G\,G_{t}^{2}-\frac{3r^{2}}{8}G_{r}G_{t}^{2}+\frac{r^{2}}{2}G\,G_{t}G_{rt}\Big)$
@@
 and
-$A_{t}$ $=2d_{r}d_{rt}+8r^{2}d_{t}d_{tt}$
+$A_{t}$ $=2d_{N,r}d_{N,rt}+8r^{2}d_{N,t}d_{N,tt}$
 $=G^{-\frac{5}{2}}\Big(-\frac{3}{32}G_{r}^{2}G_{t}+\frac{1}{8}G\,G_{r}G_{rt}-\frac{3r^{2}}{8}G_{t}^{3}+\frac{r^{2}}{2}G\,G_{t}G_{tt}\Big)$
@@
 Combining the above we arrive at
-(26) $d_{r}A_{r}+4r^{2}d_{t}A_{t}=-2\,G^{-\frac{13}{4}}\,\frac{r^{12}s^{14}(s^{2}+1)^{3}}{3s^{2}+2}.$
+(27) $d_{N,r}A_{r}+4r^{2}d_{N,t}A_{t}=-2\,G^{-\frac{13}{4}}\,\frac{r^{12}s^{14}(s^{2}+1)^{3}}{3s^{2}+2}.$
 On the other hand in $\Pi_{0}\setminus\{r=0\}$ we have, cf. (7),
-$\Delta_{{\mathbb{H}}^{n}}d$ $=d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt}$
+$\Delta_{{\mathbb{H}}^{n}}d_{N}$ $=d_{N,rr}+\frac{2n-1}{r}d_{N,r}+4r^{2}d_{N,tt}$
 $=\frac{1}{4}G^{-\frac{7}{4}}\Big\{GG_{rr}-\frac{3}{4}G_{r}^{2}+\frac{2n-1}{r}GG_{r}+4r^{2}GG_{tt}-3r^{2}G_{t}^{2}\Big\}$
@@
 From (26), (27) and (28) we obtain
-$\Delta_{p,\mathbb{H}^{n}}d=-G^{-\frac{3p+1}{4}}\,\frac{r^{3p}s^{3p+2}(s^{2}+1)^{\frac{p}{2}}\big{(}(6n+p-4)s^{2}+4n+p-5\big{)}}{3s^{2}+2}\leq 0,$
+$\Delta_{p,\mathbb{H}^{n}}d_{N}=-G^{-\frac{3p+1}{4}}\,\frac{r^{3p}s^{3p+2}(s^{2}+1)^{\frac{p}{2}}\big((6n+p-4)s^{2}+4n+p-5\big)}{3s^{2}+2}\leq 0,$
 and the desired inequality has been proved pointwise in $\Pi_{0}\setminus\{r=0\}$ (where $d_{N}$ is smooth).
 To complete the proof we argue as in the proof of Theorem 5, using in particular functions $\psi_{\epsilon}$, $\epsilon>0$, as in that proof.
-Part (i) is also used at this point since the local boundedness of $|\nabla_{H}d|$ is required when letting $\epsilon\to 0$.
+Part (i) is also used at this point since the local boundedness of $|\nabla_{H}d_{N}|$ is required when letting $\epsilon\to 0$.
 ∎
@@
 Proof.
-For simplicity we write $d$ instead of $d_{N}$. Differentiating (19) we get
-$4d^{3}d_{t}=2t-3r^{2}s+(4r^{4}s-3tr^{2})s_{t}.$
+Differentiating (20) we get
+$4d_{N}^{3}d_{N,t}=2t-3r^{2}s+(4r^{4}s-3tr^{2})s_{t}.$
 Now using the first part of (23) and the fact that $s$ solves (17) we obtain
-$d_{t}=\frac{6ts^{2}-2r^{2}s+t-9r^{2}s^{3}}{4d^{3}(3s^{2}+2)}=\frac{6ts^{2}+16r^{2}s-8t}{4d^{3}(3s^{2}+2)}.$
+(29) $d_{N,t}=\frac{6ts^{2}-2r^{2}s+t-9r^{2}s^{3}}{4d^{3}(3s^{2}+2)}=\frac{6ts^{2}+16r^{2}s-8t}{4d^{3}(3s^{2}+2)}.$
 Similarly we find that
-$d_{r}=\frac{24r^{4}s^{4}+16r^{4}s^{2}-18tr^{2}s^{3}-20tr^{2}s+6t^{2}}{4d^{3}r(3s^{2}+2)}=\frac{40r^{2}ts-32r^{4}s^{2}-12t^{2}}{4d^{3}r(3s^{2}+2)}.$
+(30) $d_{N,r}=\frac{24r^{4}s^{4}+16r^{4}s^{2}-18tr^{2}s^{3}-20tr^{2}s+6t^{2}}{4d^{3}r(3s^{2}+2)}=\frac{40r^{2}ts-32r^{4}s^{2}-12t^{2}}{4d^{3}r(3s^{2}+2)}.$
 We now let $t\to 0+$. From (17) we find
@@
 Plugging this in (20) we have
-$d^{4}(r,t)=\frac{t^{4}}{16r^{4}}-\frac{t^{6}}{64r^{8}}+O(t^{8})$
+$d_{N}^{4}(r,t)=\frac{t^{4}}{16r^{4}}-\frac{t^{6}}{64r^{8}}+O(t^{8})$
 and (i) follows. We then also have
-$\frac{1}{d^{6}(r,t)}=\frac{64r^{6}}{t^{6}}+O(\frac{1}{t^{4}})$
+$\frac{1}{d_{N}^{6}(r,t)}=\frac{64r^{6}}{t^{6}}+O(\frac{1}{t^{4}})$
 and combining the above we obtain
-$d_{t}^{2}=\frac{1}{4r^{2}}+O(t^{2})\;,\qquad d_{r}^{2}=\frac{t^{2}}{4r^{4}}+O(t^{4}).$
+$d_{N,t}^{2}=\frac{1}{4r^{2}}+O(t^{2})\;,\qquad d_{N,r}^{2}=\frac{t^{2}}{4r^{4}}+O(t^{4}).$
 We thus conclude that
-$|\nabla_{{\mathbb{H}}^{n}}d|^{2}=d_{r}^{2}+4r^{2}d_{t}^{2}=1+O(t^{2}),$
+$|\nabla_{{\mathbb{H}}^{n}}d_{N}|^{2}=d_{N,r}^{2}+4r^{2}d_{N,t}^{2}=1+O(t^{2}),$
 as required. ∎
+Remark 3.
+We note for later use that we can use (17) to eliminate $t$ from the expressions (29), (29) and (30) for $d_{N}$, $d_{N,t}$ and $d_{N,r}$ respectively. We obtain
+$d_{N}=rs(s^{2}+1)^{1/4},\qquad d_{N,r}=\frac{r^{2}s^{3}}{2d^{3}},\qquad d_{N,t}=\frac{r^{2}s^{3}}{2d^{3}}.$
 Theorem 9.
@@
 Proof.
-For simplicity we write $d$ instead of $d_{N}$. Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
+Let $E_{1},\ldots,E_{m}$ denote the sides of $\Omega$. We define
 $A_{k}=\{\xi\in\Omega:d(\xi)={\rm dist}(\xi,E_{k}),\},\qquad k=1,\ldots,m.$
@@
 We then have
-(28) $d(\xi)={\rm dist}(\xi,\partial\Pi_{k})=:d_{k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
+(31) $d_{N}(\xi)={\rm dist}_{N}(\xi,\partial\Pi_{k})=:d_{N,k}(\xi)\;,\qquad\mbox{ for all }\xi\in A_{k}.$
 It immediately follows from (31) that
-$d(\xi)=\min_{1\leq k\leq m}d_{k}(\xi),\qquad\xi\in\Omega.$
+$d_{N}(\xi)=\min_{1\leq k\leq m}d_{N,k}(\xi),\qquad\xi\in\Omega.$
 We fix a non-negative function $\phi\in{C^{\infty}_{c}}(\Omega)$ and we aim to show that
-(29) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi\,{\rm d}\xi=\sum_{k=1}^{m}\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi\,{\rm d}\xi\geq 0.$
+(32) $\int_{\Omega}\nabla_{H}d_{N}\cdot\nabla_{H}\phi\,{\rm d}\xi=\sum_{k=1}^{m}\int_{A_{k}}\nabla_{H}d_{N,k}\cdot\nabla_{H}\phi\,{\rm d}\xi\geq 0.$
 We recall the divergence theorem in the stratified setting: if the vector field $F$ takes values in the first stratum and sufficient regularity is assumed then
-[…]
+$\int_{A}{\rm div}_{H}F\,{\rm d}\xi=\int_{\partial A}F\cdot\nu_{H}dS\,$,
 where
@@
 and $\nu=(\nu_{x},\nu_{y},\nu_{t})$ denotes the usual outer normal vector.
-We cannot directly apply integration by part to each of the integrals in the RHS of (29) due to the fact that, as is seen from Lemma 5, there is a halfline $L_{k}\subset\Pi_{k}$ on which each function $d_{k}$ is not differentiable (the halfline $L_{k}$ is the image of the halfline $\{(0,0,t):t>0\}$ under the group action that maps $\Pi_{0}$ onto $\Pi_{k}$).
+We cannot directly apply integration by part to each of the integrals in the RHS of (32) due to the fact that, as is seen from Lemma 6, there is a halfline $L_{k}\subset\Pi_{k}$ on which the function $d_{N,k}$ is not differentiable (the halfline $L_{k}$ is the image of the halfline $\{(0,0,t):t>0\}$ under the group action that maps $\Pi_{0}$ onto $\Pi_{k}$).
 We define the set
 $K=\bigcup_{k=1}^{m}L_{k}$
-and so $d$ is $C^{1}$ in $\Omega\setminus K$.
+and so $d_{N}$ is $C^{1}$ in $\Omega\setminus K$.
 We now use a standard cut-off argument.
@@
 Let $k\in\{1,\ldots,m\}$. Integrating by parts and using Proposition 7 we obtain
-$\int_{A_{k}}\nabla_{H}d_{k}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi=$ $-\int_{A_{k}}\phi_{\epsilon}\,\Delta_{H}d_{k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS$
-$\geq$ $\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
+$\int_{A_{k}}\nabla_{H}d_{N,k}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi=$ $-\int_{A_{k}}\phi_{\epsilon}\,\Delta_{H}d_{N,k}\,{\rm d}\xi+\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{N,k}\cdot\nu_{k,H}\,dS$
+$\geq$ $\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{N,k}\cdot\nu_{k,H}\,dS.$
 where $\nu_{k,H}$ is defined as above relative to $A_{k}$. Adding over $k$ we arrive at
-(31) $\int_{\Omega}\nabla_{H}d\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi\geq\sum_{k=1}^{m}\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{k}\cdot\nu_{k,H}\,dS.$
+(34) $\int_{\Omega}\nabla_{H}d_{N}\cdot\nabla_{H}\phi_{\epsilon}\,{\rm d}\xi\geq\sum_{k=1}^{m}\int_{\partial A_{k}}\!\!\phi_{\epsilon}\,\nabla_{H}d_{N,k}\cdot\nu_{k,H}\,dS.$
 Now, each boundary $\partial A_{k}$ consists of outer parts where $\phi$ vanishes as well as of common boundaries with other sets $A_{j}$, $j\neq k$.
 Let us fix $k,j$ such a set $A_{k}$ and $A_{j}$ share such a common boundary $S_{kj}$.
-Letting, as above, $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ where $\nu$ is the normal vector on $S_{kj}$ which is outer with respect to $A_{k}$, we conclude that the two contributions on the surface $S_{kj}$ from $A_{k}$ and $A_{j}$ add up to
-[…]
-The surface $S_{kj}$ is a level set for the function $d_{k}-d_{j}$ and at each point $\xi\in S_{kj}$ there holds $\nabla d_{k}-\nabla d_{j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$.
+Let $\nu$ be the normal vector on $S_{kj}$ which is outer with respect to $A_{k}$ (the existence of $\nu$ is shown in the Appendix in case where the two hyperplanes are not vertical; otherwise one may use an approximation argument, see also [Lar16]).
+Defining $\nu_{H}=(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x)$ we conclude that the two contributions on the surface $S_{kj}$ from $A_{k}$ and $A_{j}$ add up to
+$\int_{S_{kj}}\!\!\phi_{\epsilon}\,(\nabla_{H}d_{N,k}-\nabla_{H}d_{N,j})\cdot\nu_{H}dS.$
+The surface $S_{kj}$ is a level set for the function $d_{N,k}-d_{N,j}$ and at each point $\xi\in S_{kj}$ there holds $\nabla d_{N,k}-\nabla d_{N,j}=\lambda\nu$ where $\lambda=\lambda(\xi)\geq 0$.
 We therefore have on $S_{kj}$
-$(\nabla_{H}\,d_{k}-\nabla_{H}\,d_{j})\cdot\nu_{H}$
-$=\big{(}\nabla_{\!x}\,d_{k}-\nabla_{\!x}\,d_{j}+2d_{k,t}y-2d_{j,t}y,\nabla_{\!y}\,d_{k}-\nabla_{\!y}\,d_{j}-2d_{k,t}x+2d_{j,t}x\big{)}$
-$\quad\quad\cdot\big{(}\nu_{x}+2y\nu_{t},\nu_{y}-2x\nu_{t}\big{)}$
+$(\nabla_{H}\,d_{N,k}-\nabla_{H}\,d_{N,j})\cdot\nu_{H}=\big(\nabla_{\!x}\,d_{N,k}-\nabla_{\!x}\,d_{N,j}+2(d_{N,k})_{t}y-2(d_{N,j})_{t}y,$
+$\quad\quad\nabla_{\!y}\,d_{N,k}-\nabla_{\!y}\,d_{N,j}-2(d_{N,k})_{t}x+2(d_{N,j})_{t}x\big)\cdot\big(\nu_{x}+2y\nu_{t},\nu_{y}-2x\nu_{t}\big)$
 $=\big(\lambda\nu_{x}+2\lambda\nu_{t}y,\lambda\nu_{y}-2\lambda\nu_{t}x\big)\cdot\big(\nu_{x}+2\nu_{t}y,\nu_{y}-2\nu_{t}x\big)$
@@
 Combining the above completes the proof of (i).
-Part (ii) is an immediate consequence of part (i) and Theorem 1.The optimality of the constant follows from Theorem 8.∎
+Part (ii) is an immediate consequence of part (i) and Theorem 1.
+The optimality of the constant follows from Theorem 9.∎
 As an immediate consequence of Theorem 11 we obtain the geometric uncertainty principle on the convex set $\Omega\subset{\mathbb{H}}^{n}$ with respect to the gauge pseudo-distance on ${\mathbb{H}}^{n}$.
@@
 Proof.
-The proof follows using standard arguments by Lemma 13, since by (42) $t\rightarrow 0$ implies that $\phi\rightarrow 0$. ∎
+The proof follows from Lemma 14 using standard arguments, since by (45) $t\rightarrow 0$ implies that $\phi\rightarrow 0$. ∎
 Theorem 16.
@@
 Proof.
-For simplicity we write $d$ instead of $d_{\rho}$.
-$({\rm i})$ We first note that by invariance under group action we may assume that $\Pi=\Pi_{0}$.
-By Lemma 13 the distance to the boundary of a point $\xi=(x,y,t)$, $(x,y)\neq(0,0)$, is given by
-$d(r,t)=\frac{\phi}{\cos\phi}r=:B(\phi)r.$
-We will see below that $d_{r}(r,t)\leq 0$. Therefore, we have
-(44) $\Delta_{{\mathbb{H}}^{n}}d=$ $\;d_{rr}+\frac{2n-1}{r}d_{r}+4r^{2}d_{tt}$
-$\leq$ $\;d_{rr}+\frac{1}{r}d_{r}+4r^{2}d_{tt}$
+$({\rm i})$ We first note that by invariance under group action we may assume that $\Pi=\Pi_{0}$. By Lemma 14 the distance to the boundary of a point $\xi=(x,y,t)$, $(x,y)\neq(0,0)$, is given by
+$d_{\rho}(r,t)=\frac{\phi}{\cos\phi}r=:B(\phi)r.$
+We will see below that $d_{\rho,r}(r,t)\leq 0$. Therefore, we have
+(47) $\Delta_{{\mathbb{H}}^{n}}d_{\rho}=$ $\;d_{\rho,rr}+\frac{2n-1}{r}d_{\rho,r}+4r^{2}d_{\rho,tt}$
+$\leq$ $\;d_{\rho,rr}+\frac{1}{r}d_{\rho,r}+4r^{2}d_{\rho,tt}$
 $=$ $\;B^{\prime\prime}(\phi)\Big(r\phi_{r}^{2}+4r^{3}\phi_{t}^{2}\Big)+B^{\prime}(\phi)\Big(r\phi_{rr}+3\phi_{r}+4r^{3}\phi_{tt}\Big)+\frac{1}{r}B(\phi).$
@@
 Substituting in (47) we arrive at
-(46) $\Delta_{{\mathbb{H}}^{n}}d\leq\frac{1}{r(Q^{\prime})^{3}}\bigg{(}4(Q^{2}+1)(B^{\prime\prime}Q^{\prime}-Q^{\prime\prime}B^{\prime})+B\,(Q^{\prime})^{3}\bigg{)}.$
+(49) $\Delta_{{\mathbb{H}}^{n}}d_{\rho}\leq\frac{1}{r(Q^{\prime})^{3}}\bigg(4(Q^{2}+1)(B^{\prime\prime}Q^{\prime}-Q^{\prime\prime}B^{\prime})+B\,(Q^{\prime})^{3}\bigg).$
 Let $A$ denote the term in large brackets in (49). We have
@@
 Therefore
-$d_{r}(r,t)=B(\phi)+B^{\prime}(\phi)\phi_{r}\,r=-\sin\phi\leq 0$
+$d_{\rho,r}(r,t)=B(\phi)+B^{\prime}(\phi)\phi_{r}\,r=-\sin\phi\leq 0$
 as claimed. Substituting we find after some computations that
@@
 Each of these four terms is non-negative for $\phi\in(0,\pi/2)$; hence $A\leq 0$ and (i) has been proved.
-(ii) The required Hardy inequality will follow by applying Theorem 1 provided we establish that $\Delta_{p,{\mathbb{H}}^{n}}d\leq 0$ in the distributional sense in $\Pi$.
-Now, by [MSC01, Theorem 3.1] (or by a direct computation for $r\neq 0$) we have $|\nabla_{{\mathbb{H}}^{n}}d|=1$ a.e.
-Hence the above condition can be simplified to $\Delta_{{\mathbb{H}}^{n}}d\leq 0$ which is in particular independent of $p>1$; see also [BFT04].
+(ii) The required Hardy inequality will follow by applying Theorem 1 provided we establish that $\Delta_{p,{\mathbb{H}}^{n}}d_{\rho}\leq 0$ in the distributional sense in $\Pi$.
+Now, by [MSC01, Theorem 3.1] (or by a direct computation for $r\neq 0$) we have $|\nabla_{{\mathbb{H}}^{n}}d_{\rho}|=1$ a.e.
+Hence the above condition can be simplified to $\Delta_{{\mathbb{H}}^{n}}d_{\rho}\leq 0$ which is in particular independent of $p>1$; see also [BFT04].
 So the result follows from (i).
@@
 Moreover $(x^{\prime},y^{\prime})$ is different from $(x,y)$ and from $(0,0)$ and there holds $d_{\rho}(\xi)<\pi r^{\prime}/2$.
-$(1b)$ Let $t>0$. The point $(0,0,t)$ has as nearest boundary points all points $(x^{\prime},y^{\prime},0)$ with $x^{\prime 2}+y^{\prime 2}=2t/\pi$.
+$(1b)$ Let $t>0$. The point $(0,0,t)$ has as nearest boundary points all points $(x^{\prime},y^{\prime},0)$ with $|x^{\prime}|^{2}+|y^{\prime}|^{2}=2t/\pi$.
 $(2a)$ Conversely, let $\xi^{\prime}=(x^{\prime},y^{\prime},0)\in\partial\Pi_{0}$ with $(x^{\prime},y^{\prime})\neq(0,0)$ and $\rho>0$ be given.
@@
 (1b) This is contained in Theorem 14.
-(2a) Assume that $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(x^{\prime},y^{\prime})$ has $\xi^{\prime}$ as its nearest boundary point and that $d(\xi)=\rho$.
+(2a) Assume that $\xi=(x,y,t)\in\Pi_{0}$ with $(x,y)\neq(x^{\prime},y^{\prime})$ has $\xi^{\prime}$ as its nearest boundary point and that $d_{\rho}(\xi)=\rho$.
 By (50) $\phi=\rho/r^{\prime}$ and the point $\xi$ is now uniquely determined by (40) and (45).
@@
 Proof.
-For simplicity we write $d$ instead of $d_{\rho}$. Let $u\in{C^{\infty}_{c}}(\Omega)$ be fixed. We consider a bounded convex polytope $\Omega^{\prime}$ such that
+Let $u\in{C^{\infty}_{c}}(\Omega)$ be fixed. We consider a bounded convex polytope $\Omega^{\prime}$ such that
 ${\rm supp}(u)\subset\subset\Omega^{\prime}\subset\subset\Omega.$
-Let us denote by $d^{\prime}$ the Carnot-Carathéodory distance to $\partial\Omega^{\prime}$. Combining Theorem 1 and Lemma 17 we obtain that
-$\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d^{\prime p}}{\rm d}\xi\;.$
-Hence, since $d^{\prime}\leq d$ in $\Omega^{\prime}$,
-$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi=\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d^{\prime p}}{\rm d}\xi\geq\Big{(}\frac{p-1}{p}\Big{)}^{p}\int_{\Omega}\frac{|u|^{p}}{d^{p}}{\rm d}\xi$
+Let us denote by $d_{\rho}^{\prime}$ the Carnot-Carathéodory distance to $\partial\Omega^{\prime}$. Combining Theorem 1 and Lemma 18 we obtain that
+$\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d_{\rho}^{\prime p}}{\rm d}\xi\;.$
+Hence, since $d_{\rho}^{\prime}\leq d_{\rho}$ in $\Omega^{\prime}$,
+$\int_{\Omega}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi=\int_{\Omega^{\prime}}|\nabla_{{\mathbb{H}}^{n}}u|^{p}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega^{\prime}}\frac{|u|^{p}}{d_{\rho}^{\prime p}}{\rm d}\xi\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|u|^{p}}{d_{\rho}^{p}}{\rm d}\xi$
 and the proof is complete. ∎
 Remark 5.
-We point that the fact that the key property that $d^{\prime}\leq d$, for $d$ and $d^{\prime}$ as in the proof of Theorem 18 in the bounded convex polytope $\Omega^{\prime}$ and reflects the geometric nature of the Carnot-Carathéodory distance, meaning in particular that that the latter is a “true” distance respecting the geometry of ${\mathbb{H}}^{n}$.
+We point that the fact that the key property that $d_{\rho}^{\prime}\leq d_{\rho}$, for $d$ and $d^{\prime}$ as in the proof of Theorem 19 in the bounded convex polytope $\Omega^{\prime}$ and reflects the geometric nature of the Carnot-Carathéodory distance, meaning in particular that that the latter is a “true” distance respecting the geometry of ${\mathbb{H}}^{n}$.
 The proof of the following corollary, which is the geometric uncertainty principle with respect to the Carnot-Carathéodory distance, follows the lines of Corollary 12 in the case of the gauge pseudodistance.
@@
 (51) $(g^{\prime}g)_{i}=\left\{\begin{array}[]{ll}g_{i}+g^{\prime}_{i}\,,&i=1,\cdots,m,\\ g_{i}+g^{\prime}_{i}+\frac{1}{2}\langle B^{(i)}g^{\prime(1)},g^{(1)}\rangle\,,&i=m+1,\cdots,n,\end{array}\right$.\,
-[…]
+where the $B^{(i)}$’s are $m\times m$ matrices, and $\langle\cdot,\cdot\rangle$ stands for the standard inner product in $\mathbb{R}^{m}$, see e.g.
+[BLU07, Remark 17.3.1].
+The group law (51) can also be written as
 (52) $g^{\prime}g=(g^{(1)}+g^{\prime(1)},g^{(2)}+g^{\prime(2)}+\frac{1}{2}\langle Bg^{\prime(1)},g^{(1)}\rangle)\,$,
-where $\langle Bg^{(1)},g^{\prime(1)}\rangle$ denotes the $(n-m)$-tuple
+where $\langle Bg^{(1)},g^{\prime(1)}\rangle$ denotes the (nm)(n-m)-tuple
 $(\langle B^{(m+1)}g^{(1)},g^{\prime(1)}\rangle,\cdots,\langle B^{(n)}g^{(1)},g^{\prime(1)}\rangle)\,.$
@@
 Example 1.
-Let $\mathbb{H}^{n}$ be the Heisenberg group and $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}$. We shall prove that the the distance $d_{N}$ (cf. (13)) is not necessarily weakly $H$-concave.
+Let $\mathbb{H}^{n}$ be the Heisenberg group and $\Pi_{0}=\{(x,y,t)\in\mathbb{H}^{n}:t>0\}$. We shall prove that the the distance $d_{N}$ (cf. (13)) is not weakly $H$-concave.
 Actually, we shall show that the weak $H$-concavity fails in a neighbourhood of any boundary point.
@@
 In this section we prove that the Euclidean distance to the boundary on a convex, bounded domain $\Omega$ is weakly $H$-concave and superharmonic.
-This provides an alternative proof of the $L^{2}$-Hardy inequality for such domains.
+This provides a proof of the $L^{p}$-Hardy inequality for such domains.
 Theorem 23.
@@
 where the last inequality follows by the choice of $v$. Hence, by the Euclidean convexity of $\Omega$, we have $h\in\Omega$, and the proof is complete. ∎
-From Theorem 22 we immediately obtain the following result; the sharpness of the constant 1/4 follows under the hypotheses of part (b) of Theorem 1.
+From Theorem 23 we obtain the following result; the sharpness of the constant follows under the hypotheses of part (b) of Theorem 1.
 Theorem 24.
-Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then
-${\rm(i)}$ $\quad\Delta_{H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
+Let $G$ be a stratified group of step two and let $\Omega\subset G$ be a bounded domain which is convex in the Euclidean sense. Then for any $p>1$ we have
+${\rm(i)}$ $\quad\Delta_{p,H}d\leq 0\;\mbox{ in the distributional sense in }\Omega;$
 ${\rm(ii)}$ The Hardy inequality
-$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{2}{\rm d}g\geq\frac{1}{4}\int_{\Omega}\frac{|\nabla_{H}d|^{2}}{d^{2}}u^{2}{\rm d}g\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
+$\hskip 42.67912pt\int_{\Omega}|\nabla_{H}u|^{p}{\rm d}g\geq\Big(\frac{p-1}{p}\Big)^{p}\int_{\Omega}\frac{|\nabla_{H}d|^{p}}{d^{p}}|u|^{p}{\rm d}g\;,\quad u\in{C^{\infty}_{c}}(\Omega),$
 is valid.
 Proof.
-By Theorem 22 the distance function is weakly $H$-concave.
-Let $X_{1},\ldots,$ $X_{m}$ be the vector fields that generate the first stratum $V_{1}$ of the corresponding Lie algebra.
-By Theorem [DGN03, Theorem 8.1] each $X_{k}^{2}d$, $k=1,\ldots,m$, is a non-positive Radon measure on $\Omega$; this proves (i).
-Part (ii) now follows from Theorem 1.
-∎
+Let $\rho\in C_{c}^{\infty}(G)$ with $\rho\geq 0$ and $\int_{G}\rho\,dx=1$. For $\epsilon>0$ we define
+$\rho_{\varepsilon}(x)=\varepsilon^{-Q}\rho(\delta_{\varepsilon^{-1}}x)\,.$
+and
+$d_{\varepsilon}(x)=(d*\rho_{\varepsilon})(x)=\int_{G}d(xy^{-1})\rho_{\varepsilon}(y)dy\,.$
+We will show that $d_{\varepsilon}$ is weakly $H$-concave on the set
+$\Omega_{\varepsilon}:=\{x\in\Omega\,:\,B_{CC}(x,\varepsilon)\subset\Omega\}\,.$
+Let $g\in\Omega_{\varepsilon}$ and let $g^{\prime}\in H_{g}\cap\Omega_{\epsilon}$. Since $g^{\prime}\in H_{g}$ there exists $v\in V_{1}$ such that
+$g^{\prime}=g\exp(v)\in\Omega_{\varepsilon},$
+and for every $\lambda\in[0,1]$ we have
+$g_{\lambda}:=g\,\delta_{\lambda}(g^{-1}g^{\prime})=g\exp(\lambda v)\,$$\qquad$
+implying that the weak $H$-concavity condition can also be written as
+$d(g\exp(\lambda v))\geq(1-\lambda)d(g)+\lambda d(g\exp v),\qquad\lambda\in[0,1]\,$,
+For $y\in\text{supp}\rho_{\varepsilon}\subset B_{CC}(e,\varepsilon)$, we define the curve
+$\gamma_{y}(t):=g\exp(tv)y^{-1},\qquad t\in[0,1]\,.$
+Then $\gamma_{y}$ is a horizontal curve inside $\Omega$ connecting $gy^{-1}\in B_{CC}(g,\varepsilon)\subset\Omega$ with $g^{\prime}y^{-1}B_{CC}(g^{\prime},\varepsilon)\subset\Omega$. So
+$d(\gamma_{y}(\lambda))\geq(1-\lambda)d(\gamma_{y}(0))+\lambda d(\gamma_{y}(1))\,$,
+that is
+$d(g\exp(\lambda v)y^{-1})\geq(1-\lambda)d(gy^{-1})+\lambda d(g\exp(v)y^{-1})\,.$
+Multiplying by $\rho_{\varepsilon}(y)\geq 0$ and integrating we obtain
+$d_{\varepsilon}(g\exp(\lambda v))\geq(1-\lambda)d_{\varepsilon}(g)+\lambda d_{\varepsilon}(g\exp v)\,$,
+i.e. $d_{\varepsilon}$ is weakly $H$-concave in $\Omega_{\varepsilon}$.
+Now, we have
+$\Delta_{p,H}d_{\varepsilon}$ $=\operatorname{div}_{H}\bigl(|\nabla_{H}d_{\varepsilon}|^{p-2}\nabla_{H}d_{\varepsilon}\bigr)$
+(56) $=|\nabla_{H}d_{\varepsilon}|^{p-4}\Big(|\nabla_{H}d_{\varepsilon}|^{2}\Delta_{H}d_{\varepsilon}+(p-2)\bigl\langle D^{2}_{H}d_{\varepsilon}\,\nabla_{H}d_{\varepsilon},\nabla_{H}d_{\varepsilon}\bigr\rangle\Big),$
+where
+$D^{2}_{H}d_{\varepsilon}:=\left(\frac{1}{2}\bigl(X_{i}X_{j}d_{\varepsilon}+X_{j}X_{i}d_{\varepsilon}\bigr)\right)_{i,j=1}^{2n}$
+is the symmetrized horizontal Hessian of $d_{\epsilon}$ (here $X_{n+k}=Y_{k}$).
+By [DGN03, Theorem 5.12] $D_{H}^{2}d_{\varepsilon}$ is negative semidefinite on $\Omega_{\varepsilon}$.
+Using also the elementary matrix inequality $A\leq{\rm tr}(A)I_{d}$ which is valid from any symmetric positive definite matrix, we conclude from (56) that
+$\Delta_{p,H}d_{\varepsilon}\leq 0\,,\qquad\text{pointwise on }\Omega_{\varepsilon}\,.$
+Finally, to show that $\Delta_{p,H}d\leq 0$ in the weak sense we need to show that for any $\varphi\in C_{c}^{\infty}(\Omega)$
+(57) $\lim_{\varepsilon\to 0}\int_{\Omega}|\nabla_{H}d_{\varepsilon}|^{p-2}\langle\nabla_{H}d_{\varepsilon},\nabla_{H}\varphi\rangle dx=\int_{\Omega}|\nabla_{H}d|^{p-2}\langle\nabla_{H}d,\nabla_{H}\varphi\rangle dx\,.$
+Since
+$\nabla_{H}d_{\varepsilon}\to\nabla_{H}d\,,\qquad\text{a.e. in supp}(\varphi)\,$,
+we have
+$\langle|\nabla_{H}d_{\varepsilon}|^{p-2}\nabla_{H}d_{\varepsilon},\nabla_{H}\varphi\rangle\to\langle|\nabla_{H}d|^{p-2}\nabla_{H}d\,\nabla_{H}\varphi\rangle$
+Now, notice that since $|\nabla_{h}d_{\varepsilon}|$ is bounded on ${\rm supp}(\varphi)$ we have
+$\left|\langle|\nabla_{H}d_{\varepsilon}|^{p-2}\nabla_{H}d_{\varepsilon},\nabla_{H}\varphi\rangle\right|\leq C_{\varphi}\,$,
+and we can apply the dominated convergence theorem to get
+$\lim_{\varepsilon\to 0}\langle\Delta_{p,H}d_{\varepsilon},\varphi\rangle=-\int_{\Omega}|\nabla_{H}d|^{p-2}\langle\nabla_{H}d,\nabla_{H}\varphi\rangle\,dg=\langle\Delta_{p,H}d,\varphi\rangle\,$,
+which completes the proof of (i) Part (ii) then follows from Theorem 1.
+Appendix
+We prove here the following result used in the proof of Theorem 11.
+Proposition 25.
+Let $\Pi_{1}$, $\Pi_{2}$ be two half-spaces in the Heisenberg group $\mathbb{H}^{n}$ and let $d_{N,1}$, $d_{N,2}$ denote the correspoding distances to the boundary with respect to the gauge quasi-norm $N$, cf.
+(13).
+If there exists $\xi\in\mathbb{H}^{n}$ such that
+[…]
+then $\Pi_{1}=\Pi_{2}$.
+Proof.
+By group action there exist $\xi_{1},\xi_{2}\in\mathbb{H}^{n}$ such that
+$d_{N,k}(\xi)={\rm dist}_{N}(\xi_{k}\,\xi,\partial\Pi_{0})=:d_{N}(\xi_{k}\,\xi),\quad k=1,2,\;\;\xi\in\Pi_{k},$
+where $\Pi_{0}$ and $d_{N}$ is as in Lemma 6.
+Given $\xi\in\Pi_{1}\cap\Pi_{2}$ we set
+$\xi_{k}\xi=:\eta_{k}=:(\alpha_{k},\beta_{k},\tau_{k})\;,\qquad k=1,2.$
+By Lemma 6 $d_{N}(\eta_{k})$ depends only on $\rho_{k}:=\sqrt{|\alpha_{k}|^{2}+|\beta_{k}|^{2}}$ and $\tau_{k}$. Let $s_{k}$, $k=1,2$, be defined by
+(58) $s_{k}^{3}+2s_{k}=\frac{\tau_{k}}{\rho_{k}^{2}}.$
+If $d_{N,1}(\xi)=d_{N,2}(\xi)$ then $d_{N}(\eta_{1})=d_{N}(\eta_{2})$ and recalling also Remark 3 we obtain
+(59) $\rho_{1}^{4}\,s_{1}^{4}\,(s_{1}^{2}+1)=\rho_{2}^{4}\,s_{2}^{4}\,(s_{2}^{2}+1).$
+Writing $\xi=(x,y,t)$ and $\xi_{k}=(x_{k},y_{k},t_{k})$ we have
+$\eta_{k}=\big(x_{k}+x,y_{k}+y,t_{k}+t+2(x_{k}\cdot y-x\cdot y_{k})\big).$
+Assume now that we additionally have $\nabla d_{N,1}(\xi)=\nabla d_{N,2}(\xi)$. There holds
+$\frac{\partial{d_{N,k}}}{\partial{t}}(\xi)=\frac{\partial{d_{N}}}{\partial{\tau_{k}}}(\eta)\,.$
+Hence, using again Remark 3,
+(60) $\rho_{1}^{2}\,s_{1}^{3}=\rho_{2}^{2}\,s_{2}^{3}.$
+Combining (59) and (60) and recalling that $s_{1}$ and $s_{2}$ are positive we obtain that
+(61) $s_{1}=s_{2}=:s\;,\qquad\qquad\rho_{1}=\rho_{2}=:\rho.$
+Now, since $d_{N}(\eta_{k})$ depends only on $\rho_{k}$ and $\tau_{k}$, we have by Remark 3 that
+$\frac{\partial{d_{N,k}}}{\partial{x_{i}}}(\xi)$ $=\frac{\partial{d_{N}}}{\partial{\rho_{k}}}\frac{\partial{\rho_{k}}}{\partial{x_{i}}}+\frac{\partial{d_{N}}}{\partial{\tau_{k}}}\frac{\partial{tau_{k}}}{\partial{x_{i}}}$
+$=\frac{1}{\rho_{k}}\frac{\partial{d_{N}}}{\partial{\rho_{k}}}(x_{i}+x_{k,i})-2\frac{\partial{d_{N}}}{\partial{\tau_{k}}}y_{k,i}$
+$=-\frac{\rho_{k}^{2}s_{k}^{4}}{d_{N,k}^{3}}(x_{i}+x_{k,i})-\frac{\rho_{k}^{2}s_{k}^{3}}{2d_{N,k}^{3}}y_{k,i},$
+and similarly
+$\frac{\partial{d_{N,k}}}{\partial{y_{i}}}(\xi)=-\frac{\rho_{k}^{2}s_{k}^{4}}{d_{N,k}^{3}}(y_{i}+y_{k,i})+\frac{\rho_{k}^{2}s_{k}^{3}}{2d_{N,k}^{3}}x_{k,i}.$
+Recalling also (61), it follows that at a point $\xi$ as in the statement we have
+$-s(x_{i}+x_{1,i})-\frac{y_{1,i}}{2}=-s(x_{i}+x_{2,i})-\frac{y_{2,i}}{2}$
+$-s(y_{i}+y_{1,i})+\frac{x_{1,i}}{2}=-s(y_{i}+y_{2,i})+\frac{x_{2,i}}{2}$
+and therefore $x_{1}=x_{2}$ and $y_{1}=y_{2}$. Finally, from (58) we also obtain $\tau_{1}=\tau_{2}$ and therefore $t_{1}=t_{2}$. Hence $\xi_{1}=\xi_{2}$ as required. ∎
+Acknowledgment. We would like to thank the two anonymous referees for the careful reading of the paper and for their useful comments.